### v2

# Unsolvable Problem Detection: Evaluating Trustworthiness of Large Multimodal Models Unsolvable Problem Detection: Robust Understanding Evaluation for Large Multimodal Models

Abstract
This paper introduces a novel task to evaluate the robust understanding capability of Large Multimodal Models (LMMs), termed Unsolvable Problem Detection (UPD). Multiple-choice question answering (MCQA) is widely used to assess the understanding capability of LMMs, but it does not guarantee that LMMs truly comprehend the answer. UPD assesses the LMM’s ability to withhold answers when encountering unsolvable problems of MCQA, verifying whether the model truly understands the answer. UPD encompasses three problems: Absent Answer Detection (AAD), Incompatible Answer Set Detection (IASD), and Incompatible Visual Question Detection (IVQD), covering unsolvable cases like answer-lacking or incompatible choices and image-question mismatches. For the evaluation, we introduce the MM-UPD Bench, a benchmark for assessing performance across various ability dimensions. Our experiments reveal that even most LMMs, which demonstrate adequate performance on existing benchmarks, struggle significantly with MM-UPD, underscoring a novel aspect of trustworthiness that current benchmarks have overlooked. A detailed analysis shows that LMMs have different bottlenecks and chain-of-thought and self-reflection improved performance for LMMs with the bottleneck in their LLM capability. We hope our insights will enhance the broader understanding and development of more reliable LMMs.
Unsolvable Problem Detection: Robust Understanding Evaluation
for Large Multimodal Models
Atsuyuki Miyai1 Jingkang Yang2 Jingyang Zhang3 Yifei Ming4 Qing Yu1,5 Go Irie6 Yixuan Li4 Hai Li3 Ziwei Liu2 Kiyoharu Aizawa1 1The University of Tokyo 2S-Lab, Nanyang Technological University 3Duke University 4University of Wisconsin-Madison 5LY Corporation 6Tokyo University of Science
1 Introduction
In recent years, following the revolutionary development of Large Language Models (LLMs) (Chen et al., 2024a; Chiang et al., 2023; Touvron et al., 2023; Wei et al., 2023), Large Multimodal Models (LMMs) (Liu et al., 2024c; Wang et al., 2023c; Hurst et al., 2024) have also demonstrated profound capabilities in various applications and significantly enhance the performance in image reasoning tasks (Antol et al., 2015; Liu et al., 2024b, e; Yue et al., 2024a).
Assessing the understanding capability of LMMs is crucial for advancing fundamental progress. Multiple-Choice Question Answering (MCQA) serves as a fundamental format for understanding evaluation and is widely used in well-established benchmarks such as MMBench Liu et al. (2024e) and MMMU Yue et al. (2024a). Each MCQA instance consists of a question paired with multiple answer options, requiring models to select the correct one. MCQA enables precise evaluation of LMMs and facilitates solid progress in the field. Consequently, many MCQA-based benchmarks have been proposed recently (Fu et al., 2024; Yue et al., 2024b; Hu et al., 2025).
Despite the advanced performance of LMMs on the accuracy of MCQA-format benchmarks, concerns remain regarding the reliability of their predictions. While previous works in the field of LLMs have discussed challenges such as maintaining invariance to different orderings of answer choices (Robinson et al., 2023; Wang et al., 2024a; Zheng et al., 2025), overcoming order sensitivity alone is not sufficient to ensure that the model truly understands the correct answer. A more recent study Wang et al. (2025) investigated LLMs’ ability to reject unsolvable problems, such as questions where the correct answer is not present among the given choices. The ability to reject unsolvable problems can serve as a more reliable means of verifying the model’s true understanding. However, this study does not focus on LMMs. When extending the evaluation from LLMs to LMMs, the types of unsolvable problems differ. Additionally, there is a lack of benchmarks and systematic evaluation protocols for comprehensively assessing recent LMMs. Consequently, existing works fail to assess the depth of LMMs’ robust comprehension.
To assess the robust comprehension of LMMs, we propose Unsolvable Problem Detection (UPD), which examines the LMM’s ability to withhold answers when faced with unsolvable problems. UPD encompasses three distinct settings: Absent Answer Detection (AAD), Incompatible Answer Set Detection (IASD), and Incompatible Visual Question Detection (IVQD). Fig. 1 shows the illustration of each setting. AAD evaluates whether the model declines to provide an answer when the correct answer is absent. IASD examines whether the model rejects a question when the given answer set is entirely incompatible. IVQD investigates the model’s ability to reject a question when there is no relevance between the image and the text question. A model that effectively rejects unsolvable problems while accurately solving standard solvable problems can be regarded as truly understanding them. On the other hand, a model that incorrectly selects an answer for unsolvable problems cannot be considered to have a true understanding of them.
For the evaluation, we introduce MM-UPD Bench, a carefully designed benchmark for evaluating UPD capability across various ability dimensions. MM-UPD employs a rigorous three-step construction process that builds upon MMBench (Liu et al., 2024e): (1) filtering out questions that can be answered by text-only language models, (2) applying the carefully designed approach for creating UPD questions, (3) finally, manually removing ambiguous samples. Built on the foundation of MMBench, our benchmarks allow us to highlight the difficulty of MM-UPD by comparing it to the self-established MMBench, and also serves as a fine-grained diagnostic tool, offering detailed insights into each LMM’s weaknesses in a broad range of MMBench’s abilities.
Our experimental results demonstrate the difficulty of MM-UPD across various state-of-the-art LMMs. The most important finding is that there is little correlation between the performance on the existing MMBench and MM-UPD Bench. This indicates that the community’s efforts to improve performance on existing benchmarks do not directly contribute to enhancing model reliability. In particular, we found that the gap between open-source and closed-source models is large, while open-source LMMs outperform closed-source LMMs on MMBench. Furthermore, our fine-grained ability analysis revealed that even closed-source models such as GPT-4o Hurst et al. (2024) exhibit weaknesses in specific abilities.
Finally, we revealed that whether the bottleneck lies in the LLM’s refusal capability or its visual understanding depends on the specific LMM. For LMMs where the bottleneck is in the LLM’s refusal capability, we observed performance improvements with LLM-driven approaches such as chain-of-thought Kojima et al. (2022) and self-reflection Kadavath et al. (2022).
Our contributions are summarized as follows:
Definition of Unsolvable Problem Detection: We propose a novel challenge called Unsolvable Problem Detection, which evaluates the LMM’s robust understanding in three problem settings: AAD, IASD, and IVQD.
Construction of MM-UPD Bench: We rigorously construct the MM-UPD Bench and provide a fine-grained diagnostic tool for broader abilities.
Benchmarking with Recent LMMs: We evaluate state-of-the-art LMMs on the UPD problem and show that our benchmarks represent a new and meaningful dimension of the performances of LMMs.
2 Related Work
Vulnerability of MCQA Evaluation. The vulnerability of MCQA has mainly been researched in the field of LLM. Previous work has aimed to mitigate bias in answer options and enhance LLMs’ consistency across different option orders (Robinson et al., 2023; Wang et al., 2024a; Zheng et al., 2025). As a more recent work, Wang et al. (2025) tested LLM’s ability to refuse unsolvable problems. They found that LLMs may perform MCQA by selecting the least incorrect option rather than distinctly correct. However, it only deals with AAD, and when applied to LMMs, the types of unsolvable problems are limited. Additionally, we consider that handling unsolvable problems requires rigorous evaluation based on ability-specific assessments, while they have not clearly identified the performance differences across abilities.
Unsolvable Problems. Unsolvable questions have been studied in NLP (Rajpurkar et al., 2018; Choi et al., 2018; Reddy et al., 2019; Sulem et al., 2022) and in VQA before the rise of LMMs (Gurari et al., 2018; Bhattacharya et al., 2019; Davis, 2020; Whitehead et al., 2022). Early VQA studies focused on task-specific models, making their benchmarks misaligned with modern LMMs due to task simplicity or differing evaluation protocols. While recent works have explored unsolvable questions in LMMs (Guo et al., 2024; Akter et al., 2024; Cao et al., 2024), they do not assess the robustness of LMMs for common MCQA.
Answer Refusal. In the task of refusing to provide an answer, there are studies in the field of LLMs that focus on abstaining due to a lack of knowledge (Kadavath et al., 2022; Feng et al., 2024). The main difference between their work and ours is that while they focus on knowledge gaps, we focus on the flaws or incompleteness of the problem itself, which leads to a different problem formulation.
3 Problem Definition
In this section, we introduce the concept of Unsolvable Problem Detection (UPD), a task designed to evaluate models’ capacity to not blindly offer incorrect answers when presented with unsolvable problems. We consider various discrepancies among the provided image, question, and answer options. Then, we categorize UPD into three distinct problem types: Absent Answer Detection (AAD), Incompatible Answer Set Detection (IASD), and Incompatible Visual Question Detection (IVQD). Here, AAD has been proposed as an unsolvable type for LLMs in existing work Wang et al. (2025), but it has not been examined with LMMs. Additionally, by incorporating IASD and IVQD, we can cover a broader scope of unsolvable types, enabling a more precise diagnosis of model weaknesses. The details of each setting are as follows:
1. Absent Answer Detection (AAD): AAD tests the model’s capability to recognize when the correct answer is absent from the provided choices. It challenges the model to not only analyze the content of questions and images but also identify when it cannot select a correct response due to the absence of an appropriate option.
2. Incompatible Answer Set Detection (IASD): IASD tests the model’s ability to identify situations where the set of answer choices is incompatible with the context.
Differing from AAD, in which the answer set is related to the question or the image, IASD deals with answer sets that are entirely irrelevant, challenging the model to withhold a response due to the lack of reasonable options. By giving a completely unrelated answer set, IASD evaluates the inherent capacity of LMMs to withhold answering, which is not affected by the granularity of the given choices.
3. Incompatible Visual Question Detection (IVQD): IVQD evaluates the LMMs’ capability to discern when a question and image are irrelevant or inappropriate. This setting tests the model’s understanding of the alignment between visual content and textual questions, aiming to spot instances where image-question pairs are incompatible.
4 Benchmarks and Evaluations
4.1 Construction of MM-UPD Bench
We create MM-UPD Bench based on MMBench (dev, 20231003) (Liu et al., 2024e). MMBench (Liu et al., 2024e) is a systematically designed benchmark for evaluating various abilities of LMMs. Utilizing MMBench allows us to assess the reliability of LMMs for general VQA questions and also enables fine-grained, ability-wise evaluation (e.g., , “Coarse Perception: Image Scene” and “Logic Reasoning: Future Prediction”).
To create MM-UPD Bench, we first filter image-agnostic questions from MMBench.
Filtering Image-Agnostic Questions. Most existing benchmarks, including MMBench, contain some image-agnostic questions (Chen et al., 2024b), which can be answered with only text information. This hinders the accurate evaluation of LMM performance. To address this issue, we first removed image-agnostic questions with text-only GPT-4 (Achiam et al., 2023). To eliminate the effect of random guessing, we applied CircularEval, which is explained in Sec. 4.4, for filtering. Next, we carefully examined the extracted question to guarantee neglectable impact of GPT-4 bias. After that, we manually eliminated the few remaining image-agnostic questions.
Next, we will construct MM-AAD, MM-IASD, and MM-IVQD, which constitute MM-UPD.
1. MM-AAD Bench: MM-AAD Bench is a dataset where the correct answer option for each question is removed.
When creating the MM-AAD Bench, we mask the correct options and remove all questions that originally have two options (which after removal would have only one option left).
To ensure no answer is present in the options, we also manually remove some questions with ambiguity.
Our MM-AAD Bench has 820 AAD questions over 18 abilities.
2. MM-IASD Bench: MM-IASD Bench is a dataset where the answer set is completely incompatible with the context specified by the question and the image.
To create MM-IASD, we shuffle all questions and answer sets and pair each question with a random answer set.
To further ensure the incompatibility, after the shuffling, we manually removed questions where the shuffled answer set was somehow compatible with the question. Our MM-IASD Bench has 919 IASD questions over 18 abilities.
3. MM-IVQD Bench: MM-IVQD Bench is a dataset where the image and question are incompatible.
This is achieved by focusing on questions that are specific, which are more likely to be incompatible with a randomly picked image.
Specifically, we first exclude the questions that can be relevant to most images (e.g., , “Which one is the correct caption of this image?”) and then shuffle the original image-question pairs.
Again, we conduct a manual check to guarantee the incompatibility of image-question pairs.
Our MM-IVQD Bench has 356 IVQD questions over 12 abilities.
In total, our UPD benchmark consists of 2,095 questions. Note here that although the MM-UPD Bench utilizes source data from MMBench, our construction approach enables us to emphasize the difficulty of MM-UPD by comparing the performance to the established MMBench, providing a deeper insight than creating an entirely new benchmark. Here, we also considered adopting MMMU Yue et al. (2024a). However, preliminary experiments showed that due to MMMU’s high difficulty level, the accuracy for standard questions was still low, making it challenging to assess reliability and potentially causing critical insights to be overlooked (as discussed in Appendix B.6). More detailed information for the construction process is provided in Appendix B.
4.2 Evaluation Metrics
To capture the ideal behavior of LMMs, we define several metrics and evaluate their performance under both standard and UPD settings. Ideal LMMs should not only yield correct answers in the standard setting (where the image, question, and answer sets are all aligned and the ground-truth answer is always within the options) but also be able to withhold answering in the UPD scenario where the question becomes unsolvable. In Fig. 2, we show the examples of these standard and UPD settings. Here, for AAD, the standard scenario refers to the correct answer included in the provided answer set. For IASD, the standard scenario refers to the correct answer included in the provided answer set and the rest options are also relevant. For IVQD, given the same question and answer set, the standard scenario has a compatible image. To better reflect the ideal behavior of LMMs, we measure several metrics throughout the paper:
1. Standard Accuracy: The accuracy on standard questions in Fig. 2.
2. UPD (AAD/IASD/IVQD) Accuracy: The accuracy of AAD/IASD/IVQD questions in Fig. 2 (AAD/IASD/IVQD).
3. Dual Accuracy: The accuracy on standard-UPD pairs, where we count success only if the model is correct on both the standard and UPD questions. This metric considers both Standard and UPD performances, making it the most suitable evaluation metric for UPD. Our evaluation thus uses this as the primary metric.
4. Original Standard:
This refers to the Standard accuracy evaluated using the prompt for the original MMBench. By adding the prompt “Answer with the option’s letter from the given choices directly" at the end of the question, it focuses specifically on improving Standard accuracy performance at the expense of UPD performance. While the Original Standard score is not Dual accuracy, we consider it the upper bound of Dual accuracy for each model based on the definition of Dual accuracy.
4.3 Evaluation Setting
To reflect the real-world use cases, we test in three settings, including a basic one and two carefully designed ones that attempt to address UPD with prompt engineering.
1. Base Setting: In the base setting, no instructions and options are provided to the model to withhold answers (shown in Fig. 2 (a)). This setting represents the most common case for using LMMs in the real world.
2. Option Setting: We add extra option “None of the above” for AAD and IASD and “The image and question are irrelevant.” for IVQD, respectively (shown in Fig. 2 (b)).
Following LLaVA (Liu et al., 2024c), we also add an instruction of “Answer with the option’s letter from the given choices directly.” to reinforce the instruction following capability.
3. Instruction Setting: We add additional instruction to explicitly gear the model towards acknowledging the unsolvable problem. The instruction is “If all the options are incorrect, answer F. None of the above.” for AAD and IASD and “If the given image is irrelevant to the question, answer F. The image and question are irrelevant.” for IVQD, respectively.
Note here that these additional options and instructions are also added to the questions in standard scenarios to make a fair comparison.
4.4 Evaluation Protocol
We adopt Circular Evaluation and GPT-involved Choice Extraction in MMBench (Liu et al., 2024e). In Circular Evaluation, a problem is tested multiple times with circularly shifted choices, and the LMM needs to succeed in all tests to pass. GPT-involved Choice Extraction first performs the matching algorithm and then uses GPT for those that do not match. To accurately identify when the model predicts as “no answer”, we leverage GPT-4o-mini (gpt-4o-mini-2024-07-18). Specifically, we count as correct for UPD questions if the model’s output is similar to “none of the above”, “I cannot answer”, or the masked correct option for AAD and IASD and “the image is irrelevant” or “I cannot answer” for IVQD. The details are shown in Appendix E.
| AAD | IASD | IVQD | ||||||||||
| Orig | Base | Opt | Inst | Orig | Base | Opt | Inst | Orig | Base | Opt | Inst | |
| Open-source LMMs | ||||||||||||
| LLaVA1.5-13b | 74.4 | 0.7 | 38.8 | 37.1 | 70.8 | 5.7 | 46.0 | 52.0 | 68.8 | 0.0 | 39.3 | 31.7 |
| LLaVA-NeXT-13B | 76.7 | 17.8 | 18.2 | 38.3 | 73.2 | 27.0 | 29.6 | 55.9 | 71.3 | 33.1 | 37.9 | 54.2 |
| LLaVA-NeXT-34B | 84.3 | 50.5 | 29.9 | 55.1 | 80.2 | 48.9 | 22.6 | 61.8 | 80.9 | 55.3 | 50.6 | 72.5 |
| LLaVA-OV-0.5B | 67.0 | 22.2 | 18.2 | 0.1 | 64.4 | 17.8 | 11.5 | 3.8 | 59.6 | 9.6 | 7.9 | 3.1 |
| LLaVA-OV-7B | 86.0 | 4.5 | 29.4 | 25.9 | 82.5 | 5.5 | 37.0 | 27.1 | 84.8 | 2.5 | 50.6 | 47.8 |
| Phi-3-Vision | 80.4 | 0.1 | 27.4 | 38.8 | 77.0 | 0.1 | 46.5 | 49.0 | 79.5 | 0.0 | 56.2 | 61.0 |
| Phi-3.5-Vision | 80.2 | 1.8 | 22.2 | 27.7 | 77.1 | 0.3 | 23.9 | 33.2 | 77.2 | 0.3 | 52.5 | 55.9 |
| CogVLM-17B | 71.5 | 0.5 | 39.3 | 3.8 | 67.7 | 0.5 | 18.3 | 4.4 | 62.9 | 0.0 | 19.4 | 9.0 |
| CogVLM2-19B | 84.0 | 0.0 | 46.1 | 44.5 | 80.8 | 0.1 | 51.6 | 58.2 | 85.4 | 0.0 | 42.7 | 42.7 |
| Idefics2-8B | 76.1 | 1.0 | 30.1 | 27.3 | 72.5 | 1.1 | 39.6 | 45.2 | 73.0 | 1.4 | 49.2 | 45.8 |
| idefics3-8B | 81.0 | 0.1 | 33.3 | 29.1 | 77.8 | 0.3 | 50.5 | 52.2 | 79.8 | 3.7 | 53.4 | 41.3 |
| InternVL2-2B | 78.2 | 6.8 | 30.6 | 17.4 | 74.2 | 14.6 | 50.6 | 17.8 | 76.4 | 15.4 | 19.9 | 14.3 |
| InternVL2-8B | 87.7 | 28.5 | 56.0 | 34.0 | 83.9 | 30.1 | 66.3 | 56.5 | 86.5 | 28.4 | 58.7 | 59.6 |
| InternVL2-40B | 91.1 | 43.5 | 55.9 | 67.9 | 87.9 | 45.0 | 59.8 | 75.7 | 90.7 | 42.7 | 56.2 | 80.6 |
| Xgen-MM | 83.2 | 0.7 | 38.3 | 31.6 | 80.0 | 0.1 | 52.1 | 42.5 | 80.9 | 0.0 | 58.1 | 35.1 |
| Qwen2-VL-7B | 84.4 | 11.5 | 38.4 | 48.3 | 81.0 | 19.7 | 49.9 | 64.0 | 80.1 | 37.1 | 63.5 | 69.1 |
| Qwen2.5-VL-7B | 88.7 | 32.2 | 49.0 | 58.5 | 84.9 | 46.1 | 70.0 | 70.4 | 84.3 | 71.1 | 74.7 | 79.5 |
| Closed-source LMMs | ||||||||||||
| GeminiPro | 72.7 | 24.5 | 40.1 | 42.9 | 70.9 | 28.1 | 48.5 | 52.1 | 69.1 | 37.6 | 57.3 | 60.4 |
| Gemini1.5Pro | 79.4 | 47.8 | 49.0 | 52.3 | 75.7 | 57.7 | 65.8 | 60.5 | 73.9 | 69.1 | 71.9 | 68.3 |
| GPT4V | 80.0 | 52.4 | 50.5 | 56.5 | 75.8 | 60.2 | 65.6 | 60.8 | 75.3 | 62.4 | 61.2 | 58.4 |
| GPT4o-mini | 78.0 | 33.5 | 48.9 | 45.1 | 75.6 | 46.5 | 63.0 | 56.9 | 72.8 | 48.3 | 58.4 | 47.5 |
| GPT4o | 83.2 | 45.6 | 57.8 | 59.3 | 80.5 | 56.1 | 68.9 | 68.0 | 76.4 | 65.2 | 69.4 | 66.0 |
| Dual | UPD | ||
| AAD | Base | 25.9 | 22.3 |
| Opt | 49.5 | 37.4 | |
| Inst | 64.9 | 22.5 | |
| IASD | Base | 27.0 | 19.6 |
| Opt | 56.5 | 42.3 | |
| Inst | 65.4 | 29.9 | |
| IVQD | Base | 14.6 | 6.5 |
| Opt | 56.7 | 35.6 | |
| Inst | 62.6 | 39.1 |
5 Experiments
5.1 Experimental Setups
We evaluated the performance of open-source and closed-source LMMs from lightweight models to 40B models. For inference, we perform a greedy search for all LMMs.
Open-source LMMs: We evaluate a range of open-source models, including InternVL2 (Chen et al., 2024c) (2B, 8B, and 40B), LLaVA series (Liu et al., 2023, 2024c, 2024d; Li et al., 2024a) (LLaVA-1.5-13B, LLaVA-NeXT-13B, LLaVA-NeXT-34B, and the latest OneVision-0.5B, 7B), Phi-3 model family (Abdin et al., 2024) (3-Vision, 3.5-Vision), CogVLM series (Wang et al., 2023c; Hong et al., 2024) (CogVLM-17B, CogVLM2-19B), Idefics series (Laurençon et al., 2024b, a) (Idefics2-8B, Idefics3-8B), Xgen-MM (Xue et al., 2024) (instruct-interleave-r-v1.5), and Qwen series (Qwen2-VL-7B (Wang et al., 2024b) and Qwen2.5-VL-7B (Team, 2025)).
Closed-source LMMs:
We evaluate GeminiPro (Team et al., 2023), Gemini 1.5 Pro (Reid et al., 2024), GPT-4V (gpt-4-vision-preview) (Achiam et al., 2023), GPT-4o mini (OpenAI, 2024), and GPT-4o (0513) (Hurst et al., 2024).
5.2 Main Results
Table 1 presents the overall Dual accuracies. Also, we show the Standard and UPD accuracies for some LMMs in Fig. 3. In Fig. 4, we show the radar charts of InternVL2-40B and GPT-4o for ability-wise fine-grained analysis.
First, we describe the two most crucial findings (F1 and F2 below).
F1: Different Performance Trends of MMBench and MM-UPD Bench.
Table 1 shows that the performance trends of MMBench (Orig) and MM-UPD (Base/Opt/Inst) are completely different. For instance, although LLaVA-OV-7B (Li et al., 2024a), CogVLM2 (Hong et al., 2024), and Xgen-MM (Xue et al., 2024) exhibit very high performance (80%) in all Original Standard, their performances in the UPD Base setting drop to less than 6% in all Base settings.
To investigate the correlation more rigorously, we calculate the correlation coefficients between the Original Standard and Dual accuracy/UPD accuracy in Table 2. We found that the correlation coefficient between UPD accuracy and the Original Standard is quite low (Max: 39.1, Min: 6.5). Dual accuracies still do not indicate a strong correlation. This suggests that our benchmark is capable of accurately capturing an important aspect of trustworthiness that has not been measured by previous benchmarks.
F2: Large Gap between Open-source LMMs and Closed-source LMMs.
As shown in Table 1, there is a significant performance gap between open-source LMMs and closed-source LMMs. One of the reasons for this performance gap is the training difference: Closed-source models are trained for refusal considering real-world user applications according to their system cards (Hurst et al., 2024; OpenAI, 2023). On the other hand, open-source models usually compete for performance with limited publicly available benchmarks.
However, among open-source LMMs, models with large LLMs such as LLaVA-NeXT-34B and InternVL2-40B demonstrate performance comparable to closed-source models. Compared to smaller models trained on the same VQA data, such as LLaVA-NeXT-13B and InternVL2-2B/8B, there is a significant performance improvement, suggesting that the performance of the base LLM also plays a crucial role.
However, a detailed check of each output reveals that a quality gap still exists between these powerful open-source LMMs and closed-source LMMs (refer to Appendix F.2).
InternVL2-8B | GPT-4o | |||||
| Base | 17.8 (72.6/23.2) | 4.5 (85.4/5.1) | 28.5 (82.7/30.2) | 45.6 (80.2/52.3) | ||||
| AAD | CoT | 42.8 (60.0/60.5) | 37.9 (77.1/42.8) | 29.0 (83.7/29.6) | 47.7 (77.9/56.0) | |||
| Self-reflection | 37.8 (66.2/50.0) | 27.6 (84.6/29.1) | 38.7 (81.5/41.2) | 55.2 (69.8/75.1) | ||||
| Base | 27.0 (68.9/40.8) | 5.5 (81.8/5.7) | 30.1 (78.3/35.0) | 56.1 (77.9/70.0) | ||||
| IASD | CoT | 43.9 (56.4/70.8) | 36.7 (73.7/45.7) | 29.4 (79.5/32.5) | 48.4 (74.5/64.2) | |||
| Self-reflection | 36.7 (62.6/55.8) | 35.4 (81.1/45.2) | 34.0 (77.4/41.0) | 57.9 (61.8/83.6) | ||||
| Base | 33.1 (67.4/44.9) | 2.5 (85.4/3.1) | 28.4 (82.3/35.1) | 65.2 (73.6/90.2) | ||||
| IVQD | CoT | 47.5 (59.0/75.3) | 14.9 (75.3/18.0) | 14.9 (83.1/17.1) | 57.2 (70.5/83.4) | |||
| Self-reflection | 39.0 (59.8/61.5) | 31.7 (85.4/34.6) | 30.3 (81.2/37.9) | 57.9 (61.8/96.1) |
Next, we provide detailed findings below to support the rationale behind the above findings.
F3: UPD Score is Significantly Lower than Standard in Base and Solution Varies by LMMs.
Fig. 3 shows the Standard (blue) and UPD (red) accuracy. The performance was compared, with each row showing the results for AAD, IASD, and IVQD, and each column showing the results for Base, Option, and Instruction. Models (i)-(v) in the figure denotes open-source models and Model (vi)-(viii) denotes closed-source models.
First, for the Base settings, open-source LMMs indeed exhibit lower UPD accuracy compared to Standard accuracy. Even for the Option setting, open-source LMMs still tend to perform worse on UPD than on Standard. When additional instruction is added, some models finally show a reversal in UPD and Standard performance. However, for (i) LLaVA-OV-7B and (iii) InternVL2-8B, the UPD accuracy decreases compared to the Option setting.
Therefore, effective prompting strategies to refrain from providing answers vary by LMMs.
F4: Performance of AAD, IASD, and IVQD Diagnose Each LMM’s Weakness.
The weaknesses of each model can be diagnosed by examining the performance differences in AAD, IASD, and IVQD. Regarding IVQD, even in Base settings, closed-source models demonstrate high UPD performance (Fig. 3 (vi)-(viii) in IVQD), whereas open-source models show significantly lower UPD performance (Fig. 3 (i)-(v) in IVQD).
In the comparison between AAD and IASD, models such as LLaVA-OV-7B and Phi3.5V exhibit low UPD accuracy under both Base settings (Fig. 3 (i)-(ii) in AAD and IASD), indicating that these models inherently lack the refusal ability, regardless of the option’s semantics.
On the other hand, other LMMs show high UPD performance in IASD Base setting while they have difficulty for AAD Base setting (Fig. 3 (iii)-(viii) in AAD and IASD), which indicates they possess a certain level of refusal capability, but the option’s granularity affects the performances a lot.
F5: Performance Trends Vary across Abilities.
Fig. 4 presents the detailed scores for each ability of InternVL2-40B and GPT-4o. These results reveal that the ease of withholding responses varies by ability.
Thus, by examining the ability-wise scores, we can more clearly identify each model’s weaknesses.
5.3 Analysis
To determine whether the issue lies with the vision or language side, we tested if the LMM could correctly choose “None of the above" when directly given the answer in the prompt. For example, we prompted: “$Question (How many cows are…) The answer is three. Choose the option that best fits the above answer. A. two B. four C. eight D. None of the above." If the LMM answers correctly, the issue likely stems from unstable image understanding; if not, it is a limitation of the LLM.
The experimental results are shown in Fig. 5. GPT-4o was found to successfully refuse in most abilities and the next challenge lies in improving image understanding. While InternVL2 does not match GPT-4o, it has relatively high performance, highlighting that improving image understanding is a future challenge. On the other hand, it was found that LLaVA-NeXT-13B, LLaVA-OV, and Qwen2VL have very low performance on the language side itself (fine-tuned Vicuna1.5-13B (Chiang et al., 2023) for LLaVA-NeXT-13B, and fine-tuned Qwen2-7B (Yang et al., 2024a) for LLaVA-OV and Qwen2VL).
Based on these results, we hypothesized that for models with a bottleneck on the language side, approaches aimed at improving language capabilities, such as chain of thought Kojima et al. (2022) and self-reflection Kadavath et al. (2022), would be effective. The results of these approaches are presented in Table 3. As expected, we found that these approaches were indeed effective for models with a bottleneck on the language side, such as LLaVA-OV and LLaVA-NeXT. We also examine the performance of instruction tuning. The detail of these additional experiments is included in Appendix C.2.
6 Conclusion
This paper proposes the UPD challenges for LMMs. For the UPD challenge, we introduce the MM-UPD Bench. Our experimental results indicate the difficulty of MM-UPD across various state-of-the-art LMMs and reveal a new aspect of reliability that cannot be measured by existing benchmarks.
Limitation
Proposing Innovative Approach for UPD. This study primarily focuses on the rigorous task design of UPD and proposing approaches is left as an important future work. We applied existing methods and crucial baseline approaches, clarifying the efficacy and limitations of each method. Building on our findings, to develop novel methods will be an important future work.
Extension to More Diverse Questions.
MM-UPD Bench provides general multiple-choice QA datasets. We did not add more challenging questions, as the current models still struggle with standard questions (refer to Appendix B.6). However, as LMMs advance, incorporating these difficult questions into UPD will be an important future work.
Appendix
Appendix A Additional Related Work
Large Multimodal Model (LMM). Recent advancements in multimodal models have been driven by innovative training methods (Chen et al., 2020; Zhou et al., 2020; Zhang et al., 2021; Li et al., 2020; Alayrac et al., 2022; Awadalla et al., 2023). Following the success of large language models (LLMs), many LMMs have been developed with improved instruction-following capabilities (Liu et al., 2023, 2024c, 2024d; Li et al., 2024a; Dai et al., 2023; Zhu et al., 2024; Zhang et al., 2024c; Gao et al., 2023; Ye et al., 2023, 2024; Zhao et al., 2023; Li et al., 2023a; Monajatipoor et al., 2024; Zhao et al., 2024; Li et al., 2025; Lin et al., 2024; Zhang et al., 2024b). Additionally, closed-source LMMs like GPT-4V (Achiam et al., 2023), GPT-4o (Hurst et al., 2024), and Gemini (Team et al., 2023) have exhibited strong performance across various vision-language tasks. However, a significant challenge remains in accurately evaluating the trustworthiness of these LMMs, highlighting the need for more robust and comprehensive benchmarks.
LMM Benchmarks. As multi-modal pretraining and instruction tuning has gained prominence, the previous standard evaluation benchmarks e.g., VQA (Antol et al., 2015; Goyal et al., 2017), OK-VQA (Marino et al., 2019), COCO (Lin et al., 2014), and GQA (Hudson and Manning, 2019) become insufficient (Yue et al., 2024a, b). To more comprehensively assess the capabilities of LMMs, recent efforts have introduced benchmarks such as SEED (Li et al., 2024b), LLaVA-Bench (Liu et al., 2023), MMBench (Liu et al., 2024e), MM-Vet (Yu et al., 2024), MathVista (Lu et al., 2024), Mathverse (Zhang et al., 2024d), MMStar (Chen et al., 2024b), BLINK (Fu et al., 2024), MMMU (Yue et al., 2024a), and MMMU-Pro (Yue et al., 2024b) have emerged and become common benchmarks for evaluating LMMs (Li et al., 2024a). Among these, MMBench provides evaluations across a broad range of fine-grained abilities, which is highly important for assessing UPD. Therefore, by adopting MMBench, we can (i) evaluate performance across a wider range of tasks compared to similar recent works (Guo et al., 2024; Akter et al., 2024; Cao et al., 2024) that adopt conventional benchmarks (Lin et al., 2014; Goyal et al., 2017), and (ii) emphasize the challenge of UPD by comparing standard MMBench performance with UPD performance.
Model Hallucinations. In LMMs, “hallucination” typically refers to situations where the generated responses contain information that is inconsistent in the visual content (Rohrbach et al., 2018; Wang et al., 2023b; Zhou et al., 2024; Guan et al., 2024; Sun et al., 2024; Cui et al., 2023; Jiang et al., 2024). Recent LMMs, such as LLaVA (Chung et al., 2024; Liu et al., 2024c), have also encountered the challenge of hallucination (Jiang et al., 2024). To evaluate hallucination in LMMs, various benchmarks, POPE (Li et al., 2023b), M-HalDetect (Gunjal et al., 2024), GAVIE (Liu et al., 2024a), HallusionBench (Guan et al., 2024), and Bingo (Cui et al., 2023) have been proposed. Hallucination evaluation and detection (Li et al., 2023b; Wang et al., 2023b; Liu et al., 2024a), and hallucination mitigation (Yin et al., 2024; Zhou et al., 2024; Gunjal et al., 2024; Liu et al., 2024a; Favero et al., 2024; Huang et al., 2024; Park et al., 2024; Wang et al., 2024c) have also been explored. These existing studies deal with a wide range of hallucination issues. Unlike previous works, we address the hallucination issues where the LMM produces incorrect responses when presented with unsolvable problems. Only a few very recent works have addressed this type of hallucination (Guo et al., 2024; Akter et al., 2024; Cao et al., 2024). However, they do not assess the robustness of LMMs for common MCQA.
AI Safety. A reliable visual recognition system should not only produce accurate predictions on known context but also detect unknown examples (Amodei et al., 2016; Mohseni et al., 2022; Hendrycks et al., 2021; Hendrycks and Mazeika, 2022). The representative research field to address this safety aspect is out-of-distribution (OOD) detection (Hendrycks and Gimpel, 2017; Liang et al., 2018; Yang et al., 2024b, 2022; Zhang et al., 2024a). OOD detection is the task of detecting unknown samples during inference to ensure the safety of the in-distribution (ID) classifiers. Along with the evolution of the close-set classifiers, the target tasks for OOD detection have evolved from the detectors for conventional single-modal classifiers to recent CLIP-based methods (Miyai et al., 2024; Hendrycks and Gimpel, 2017; Yu and Aizawa, 2019; Wang et al., 2021; Du et al., 2022; Ming et al., 2022b; Esmaeilpour et al., 2022; Ming et al., 2022a; Yang et al., 2023; Wang et al., 2023a; Miyai et al., 2023a, b). The next crucial challenge is to evolve the problems faced in OOD detection to LMMs in the VQA task. We consider that our UPD is an extension of the concept of OOD detection, where the model should detect and not predict unexpected input data.
Appendix B Benchmark Construction
We carefully adapt MMBench (validation) to create our MM-UPD Bench. For simplicity of explanation, we show the mapping table of each index and each ability in MMBench in Table A. MMBench (20231003) is a VQA dataset consisting of 1,164 questions. To create the MM-UPD Bench from MMBench, we conduct the following processes.
B.1 Processing for MMBench Adaptation
First, we performed the following steps for the original MMBench to ensure the quality of our benchmarks.
Exclusion of Image-Agnostic Questions. In the original MMBench, a subset of the questions were image-agnostic questions, which can be answered with only text information. To ensure the validity of the LMM benchmark, we carefully excluded these questions. First, we removed the questions that could be accurately answered by text-only GPT-4. To eliminate the effect of random guessing, we applied CircularEval for filtering. This process extracted 124 questions as image-agnostic questions. To investigate GPT-based biases, we thoroughly examined all the 124 questions excluded by GPT-4. As a result, we found that 110 of 124 were questions that could be answered using only the question texts. The remaining 14 questions appeared image-specific but could be answered by GPT-4 using information from its training, such as the frequency of words in the answer options. However, these 14 questions were primarily limited to common questions in the benchmark. Therefore, the impact of removing these 14 questions is considered to be minimal and we have confirmed that our filtering process does not introduce bias from GPT-4. Then, we manually checked and excluded the few remaining image-agnostic questions. In total, we removed 13% of the original questions as image-agnostic questions. Therefore, we argue that our benchmark consists of image-dependent questions.
Exclusion of Image Quality Ability. In the original MMBench, the Image Quality ability questions consist of 31 two-choice questions and 22 four-choice questions. We removed the two-choice questions in the AAD settings so that more than one choice remains after masking the choices. As for the remaining four-choice questions in Image Quality, our preliminary experiments indicated that these questions proved to be extremely difficult even with the original standard settings. Since it is difficult to measure accurate UPD performances with the questions that is extremely difficult even for the Standard setting, we removed the Image Quality ability.
Exclusion of Options related “None of the above". We remove the questions that originally had options related “None of the above" in order to guarantee that no correct option exists after masking the correct option. Specifically, a few questions have the option of “None of these options are correct." or “All above are not right". Since these options are not correct answers for the original questions, we simply deleted such options.
Clarification of the Semantics of the Options. We clarify the meaning of the options. Specifically, some questions in #6: Attribute Comparison have “Can’t judge”. “Can’t judge” means that “I can’t judge from the image since the image does not have enough information”. However, “Can’t judge” might be interpreted as “Since the given options are incorrect, can’t judge.” Therefore, we changed the option of “Can’t judge” to “Can’t judge from the image due to the lack of image information” to reduce the ambiguity.
After the above adaptation process, we construct MM-UPD Bench (MM-AAD, MM-IASD, MM-IVQD) as follows:
B.2 Construction of MM-AAD Bench
When creating the MM-AAD Bench, we mask the correct options and remove all questions that originally have two options (which after removal would have only one option left). Also, we remove the questions whose answer is “both A,B, and C" and “all of these options are correct". To ensure no answer is present in the options, we also manually remove some questions with ambiguity where one of the remaining options is very similar to the masked correct option (e.g., Q. What can be the relationship of these people in this image? Masked Option: Friends, Similar remaining option: Colleagues). Our MM-AAD Bench has 820 AAD questions over 18 abilities. The distribution of questions for each ability is shown at the top of Table B.
B.3 Construction of MM-IASD Bench
To create MM-IASD, we shuffle all questions and answer sets and pair each question with a random answer set. To further ensure the incompatibility, after the shuffling, we manually removed questions where the shuffled answer set was somehow compatible with the question (e.g., Q. Which of the following captions best describes this image? Correct answer: A person holding a bouquet of flowers, Similar shuffled option: Happiness). Our MM-IASD Bench has 919 IASD questions over 18 abilities. The distribution of questions for each ability is shown in the middle of Table B.
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | ||||||||||||
| OCR |
| #8 | #9 | #10 | #11 | #12 | #13 | ||||||||||||||
| #14 | #15 | #16 | #17 | #18 | ||||||||||
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #10 | #11 | #12 | #13 | #14 | #15 | #16 | #17 | #18 | total | |
| AAD | 35 | 94 | 62 | 50 | 49 | 44 | 45 | 15 | 32 | 38 | 46 | 29 | 44 | 25 | 31 | 42 | 93 | 46 | 820 |
| IASD | 39 | 97 | 77 | 54 | 53 | 39 | 43 | 20 | 42 | 41 | 63 | 42 | 43 | 35 | 33 | 49 | 98 | 51 | 919 |
| IVQD | 31 | 68 | 36 | 18 | 14 | 23 | 45 | 15 | 43 | - | 16 | 23 | - | - | - | - | 24 | - | 356 |
| Ability | Example of removed question | ||
| #3 Object Localization | How many dogs are in this picture? | ||
| #15 Image Topic | Which one is the correct caption of this image? | ||
| #16 Image Emotion | Which mood does this image convey? | ||
Which Python code can generate the content of the image? | ||
| #14 Future Prediction | What will happen next? | ||
| #10 Identity Reasoning | What’s the profession of the people in this picture? | ||
| #18 Image Style | Which style is represented in this image? |
B.4 Construction of MM-IVQD Bench
To create MM-IVQD Bench, we first exclude the questions that can be relevant to most images and then shuffle the original image-question pairs. In Table C, we show some representative examples of removed questions. For example, the question of “How many …" can be compatible with any image, since the correct option of “None of the above" always exists for any image even when the image has no corresponding objects. For the question of “What’s the profession …", we can interpret the profession from any kind of image (e.g., A beautifully captured image would suggest the profession of a photographer). In addition, we exclude the option “Can’t judge from the image due to the lack of image information.” because this option can be a correct answer for IVQD questions. Again, we conduct a manual check to guarantee the incompatibility of image-question pairs. Our MM-IVQD Bench has 356 IVQD questions over 12 abilities. The distribution of questions for each ability is shown in the bottom of Table B. Here, the lack of some ability (e.g.,#16 Image Emotion) indicates that there are many removed questions that can be applied to any image. Note that the small number of IVQD questions compared to AAD and IASD is due to our careful annotation and that even this number of questions is sufficient to show the performance difference between each LMM and method from our main experimental results.
Here, one might wonder why we exclude questions rather than modify them. That is true that we can increase the number of questions by making the general question more specific. However, these question types are inherently less likely to encounter IVQD situations, and there is a concern that forcibly modifying the questions might lead to a divergence from real-world IVQD distribution. Moreover, incorporating numerous question types with low IVQD frequency could overshadow the significance of question types that are more likely to occur, thereby compromising the accurate assessment of IVQD performance. Therefore, we chose to exclude these questions rather than modify them.
B.5 Manual Curation Procedure
The dataset curation is carried out by four annotators from the authors. To improve the efficiency of collaborative curation and ensure consistency in quality, we first transcribed the image-question pairs from MMBench into an online editing tool (i.e., Google Docs) and conducted the curation process directly within the platform. To enhance the consistency, each question was independently reviewed by two annotators. Finally, the lead author verified the validity of all curation. If a problem needed to be refined, the reason was recorded in detail as a comment. For example, in the case of IVQD, which required the most careful curation, one annotator would leave a comment on points such as “The reason the image relates to the question is…" or “If we change this image into …, the irrelevance is guaranteed.". If another annotator agreed with the comment, the problem was refined. In cases where the other annotator disagreed, all four annotators engaged in discussions to reach a consensus.
We consider that collaborative tools such as Google Docs, double-checking by two annotators, and detailed justifications with collective decisions ensure curation consistency.
B.6 Validity of UPD Benchmark on More Complex Datasets
The reason for the exclusion of the recent challenging dataset (e.g., MMMU (Yue et al., 2024a)) for our UPD benchmark is that the evaluation significantly deviates from the aspect of reliability and potentially causes us to miss important findings. To verify this, we conducted experiments with MMMU in the AAD setting.
Setup. As preprocessing, we first removed about 24.2% of image-agnostic questions from the MMMU’s validation set (900 questions) using GPT-4-based CircularEval. Then, to improve the interpretability of scores, we utilized only multiple-choice questions with four options (which make up the majority of questions in MMMU) and created MMMU-AAD using the same pipeline of MM-UPD. MMMU-AAD consists of 459 questions. For the evaluation of MMMU-AAD, we applied the CircularEval strategy as used in MM-UPD.
| Orig. | Base | Opt | Inst | |
| LLaVA-OV-7B | 23.5 | 0.7 (20.5, 5.7) | 0.7 (22.4/2.4) | 0.7 (20.0/2.4) |
| InternVL2-8B | 24.4 | 4.1 (19.8, 9.4) | 2.8 (22.0, 4.1) | 3.5 (21.8, 11.8) |
| LLaVA-NeXT-34 | 23.9 | 6.3 (12.0, 35.4) | 0.4 (23.4, 1.8) | 4.2 (9.6, 59.7) |
| GPT-4o | 27.5∗ | 15.5 (42.9, 20.9) | 8.9 (24.4, 19.0) | 23.7 (35.9, 48.4) |
Result. We show the comparison results in Table D. Based on these results, in contrast to MM-UPD, we could not verify the efficacy of either the Option or Instruction approaches. This result reveals that the evaluation using MMMU fails to capture important findings of the effectiveness of these prompting approaches for UPD. Specifically, for expert-level problems, LMMs do not have accurate answers due to the lack of capability. Therefore, even if they choose an incorrect option when encountering an unsolvable problem, this only indicates a lack of reasoning ability or knowledge and does not necessarily demonstrate a lack of refusal ability. Additionally, due to the very low overall performance, it becomes difficult to have meaningful discussions based on these minute differences in scores. Therefore, we exclude datasets with low Standard accuracy.
Appendix C Experimental Detail
C.1 Experimental Setup
Computing Infrastructures. We conduct all our evaluations of open-source models on a single NVIDIA A100 (80GB) GPU.
HyperParameters of LMM Inference. We set a temperature to 0 for all models during inference.
Base | Opt | Inst |
| AAD | 76.7 | 68.9 | 18.3 | 18.2 | 38.8 | 47.6 | ||||||
| IASD | 73.2 | 65.4 | 31.4 | 29.8 | 57.8 | 60.0 | ||||||
| IVQD | 71.3 | 67.4 | 29.8 | 37.9 | 54.2 | 59.6 |
Base | Opt | Inst |
| AAD | 84.3 | 78.6 | 53.2 | 29.9 | 55.2 | 63.8 | ||||||
| IASD | 80.2 | 74.8 | 56.7 | 22.6 | 61.9 | 73.3 | ||||||
| IVQD | 80.9 | 74.7 | 53.4 | 50.6 | 72.5 | 70.2 |
C.2 Detail of LLM-driven Methods
In this section, we explain the details of the LLM-driven approaches in Sec. 5.3.
Chain of Thought (CoT) Prompting. In this experiment, we investigate whether a widely used Zero-shot CoT (Kojima et al., 2022) is effective for UPD. We added the prompt “Let’s think step by step." at the end of the prompt and measured the performance.
Self-reflection Self-reflection is a method that allows the model to reflect on its own responses (Kadavath et al., 2022). It has been shown that LLMs might have preliminary capabilities for judging and evaluating their own answers (Kadavath et al., 2022; Feng et al., 2024). In this experiment, we evaluate whether self-reflection is effective for UPD. We show the prompt for self-reflection in Table F. We prompt the LMM to self-reflect directly after its generated answer with the phrase “The above answer is: 1. True 2. False,” following LLM protocols (Kadavath et al., 2022; Feng et al., 2024). For evaluation, if the LMM outputs “2. False," the response will be withdrawn. Otherwise, we use the original LMM’s response for the evaluation.
Appendix D Additional Experiments
We explore effective instruction-tuning recipes for solving UPD. To solve all kinds of UPD problems, we meticulously designed the data distribution for instruction tuning on Standard, AAD, IASD, and IVQD questions.
D.1 Setup
Dataset. For the dataset, we use a subset of an open-knowledge VQA dataset, A-OKVQA (Schwenk et al., 2022). It is a single-choice type VQA dataset that has been used for training InstructBLIP (Dai et al., 2023) and LLaVA-1.5 (Liu et al., 2024c). The samples in A-OKVQA do not overlap with our benchmarks.
To address all three types of problems, the ratio of the tuning data for each task is important. Therefore, we examine the difficulty and heterogeneity of each task and then seek the optimal amount and proportion of each type of question. We first create 4 kinds of datasets for standard questions, AAD questions, IASD questions, and IVQD questions, respectively. For each dataset, we include the questions for the base setting and the questions with additional options. For AAD/IASD/IVQD datasets, we set “I cannot answer.” as the answer for the base-setting questions and set the UPD-specific options such as “None of the above” to the answer for the option-setting questions. Also, to make it robust for the number of options, we create the questions with 2-4 options by augmentations.
Model and Tuning Method. The experiments were conducted based on LLaVA-NeXT-13B/34B due to its ease of implementation and its powerful performance. We adopt LoRA tuning (Hu et al., 2022) by considering the effectiveness and low memory usage.
D.2 Analysis
In this section, we aim to explore the optimal tuning recipe. First, we investigate the difficulty and heterogeneity of the AAD, IASD, and IVQD tasks. Then, by conducting experiments with varying proportions of each task and adjusting the amount of data, we identify the best tuning recipe.
Difficulty and Heterogeneity of Each Task. To create a dataset that addresses all UPD problems, it is crucial to examine the difficulty and heterogeneity of each task. To this end, we compare the performances when we use only one UPD dataset from all three kinds of UPD datasets, which indicates the difficulty or similarity of each task. In Table 7(b), we show the result. From this result, we find that, for AAD and IVQD, we need to include their own training data, while both IVQD and AAD data are sufficient to solve IASD questions. This is because IASD can be considered a simpler version of the AAD question since the answer-set does not include the correct answer, and it is also related to IVQD since the answer-set is not related to the given image. Hence, to reduce the complexity, we can create the tuning dataset from AAD and IVQD data.
Ablation on Ratio of Each UPD Task. In Fig. C, we illustrate the relationship between the ratio of Standard, AAD, and IVQD instruction tuning data and the performance of each UPD, Standard, and Dual accuracy. We set the ratio of Standard: AAD: IVQD to 3.3:3.3:3.3, 6:2:2, 7:2:1, 1:0:0. From this result, increasing the ratio of UPD tuning data, the UPD performance improved much while the standard accuracy degrades. Conversely, increasing the proportion of Standard data degrades the UPD performance. We can see that the ratio of 6:2:2 is an effective ratio for all the settings.
Ablation on Data Size. In Fig. C, we illustrate the relationship between the tuning data size and the performance of each UPD, Standard, and Dual accuracy. In this experiment, we set the ratio of Standard, AAD, and IVQD is 0.6, 0.2, and 0.2. From this result, 10,000 samples are enough to tune for our LoRA-based instruction tuning.
From these experiments, we find that the most effective approach is to include 20% AAD and 20% IVQD questions each, and 10,000 samples are sufficient for tuning.
D.3 Result
Table E demonstrates that instruction tuning is effective for UPD, showing the performance efficacy and limitations with UPD-specific training. However, UPD-specific training may degrade the performance of other general tasks. Therefore, if the user intends to use LMMs for broader, more general purposes rather than just for UPD tasks, instruction tuning may not be a good approach. It is a future challenge to propose a method that improves UPD performance while maintaining performance on general tasks.
| ${Question} |
| Your Previous Answer: <LMM’s Answer> |
| The above answer is: |
| 1. True |
| 2. False |
| Answer with the letter of either option: 1 or 2 directly. |
| Training Data | AAD | IASD | IVQD |
| Standard+AAD | 66.5 | 72.9 | 51.7 |
| Standard+IASD | 45.2 | 74.4 | 26.7 |
| Standard+IVQD | 52.1 | 72.2 | 73.6 |
| Training Data | AAD | IASD | IVQD |
| Standard+AAD | 73.9 | 96.4 | 63.8 |
| Standard+IASD | 46.7 | 96.1 | 32.0 |
| Standard+IVQD | 55.8 | 94.7 | 95.8 |
Appendix E Evaluation
E.1 Further Discussion of Evaluation Metrics
We consider the Original Conditional Dual accuracy (OC-Dual) score, a metric that takes into account the Original Standard Accuracy for each LMM. Dual Accuracy is an evaluation metric that equally assesses Standard accuracy and UPD accuracy. This metric inherits the widely supported concept of a reliable model that answers when it should and refuses when it should not (Amodei et al., 2016; Hendrycks et al., 2021; Yang et al., 2024b). However, it also takes into account differences in the original capability for Standard problems. Therefore, we consider the OC-Dual score as a score that does not depend on the original capability. The OC-Dual score is defined as follows: OC-Dual = (Success in all Original Standard, Standard, UPD settings) / (Success in Original Standard).
We plotted the relationship between OC-Dual accuracy and Dual accuracy in Fig A. To quantify the relationship between these scores, we calculated the correlation coefficient () and Spearman’s rank correlation coefficient (). The analysis revealed a very strong correlation between the two metrics. This is attributed to the fact that the Original Standard performance of current LMMs shows little variation within the MM-UPD Bench. Given that OC-Dual accuracy does not guarantee practical usability, the Dual accuracy for MM-UPD is the most effective to precisely assess the reliability of state-of-the-art LMMs without compromising real-world applicability.
E.2 Automatic Evaluation Strategy
We adopt Circular Evaluation and GPT-involved Choice Extraction in MMBench (Liu et al., 2024e) as an evaluation strategy. In Circular Evaluation, a problem is tested multiple times with circularly shifted choices, and the LMM needs to succeed in all testing passes. GPT-involved Choice Extraction first performs the matching algorithm and then uses GPT for those that do not match.
However, since the existing MMBench evaluations are optimized for standard questions, directly using them would assign standard options to refusal responses. Therefore, we made the following modifications for the UPD challenge.
Simplification of the Matching Algorithm. To apply the matching algorithm for UPD, we simplify the matching algorithm to prevent the refusal responses from matching the given options. In detail, when an option is denoted simply by a letter such as ‘A’ or expressed as ‘A) XXX’, ‘A. XXX’, ‘A, XXX’, ‘(A) XXX’ without the inclusion of other choices within the ‘XXX’ portion, it is considered that ‘A’ is being predicted.
Change of the Template for GPT Evaluation. Next, to identify the refusal prediction, we leverage GPT following MMBench. We leverage GPT-4o-mini (gpt-4o-mini-2024-07-18), considering its high performance and low cost.
We slightly change the template for the original MMBench, and create the query template for each setting in Fig. D. As for ${option}, we add UPD-specific options to recognize UPD predictions. In Fig. E, we illustrate the options for each setting. For AAD, we add two options: a masked correct option, and the option of “The correct answer is No answer, None of the above, all provided options are incorrect, or I cannot answer.”. For IASD, we add two options: a masked correct option, and the option of “The correct answer is No answer, None of the above, all provided options are irrelevant or incorrect, or I cannot answer.”. For IVQD, we add an option of “The correct answer is that The image is incompatible with the question, or I cannot answer.” For the additional-instruction setting, we also add the option “F. None of the above” or “F. The image and question are irrelevant.”. In each setting, we regard the options indicated by check marks (Fig. E), as correct ones.
E.3 Comparison to Human Decision
In Fig. F, we investigate the alignment of scores given by GPT-4o-mini and humans for the base setting. To investigate the performance of the UPD predictions, we sampled every 100 predictions of LLaVA-Next-34B and GPT-4o output that were not matched by pattern matching and manually evaluated them. We found that the match rate with human evaluations is sufficiently high.
Appendix F Error Analysis
F.1 Failure Examples of GPT-4o
We show some GPT-4o’s failure examples in Fig G, H, and I. GPT-4o is weak in the following categories in AAD: #3: Object Localization, #6: Attribute Comparison, #7: Nature Relation, and #12: Physical Property Reasoning, so we included examples of these abilities. From this result, it is clear that it selects answers from incorrect options.
There are two interesting discoveries. The first point is that GPT-4o tends to select the option that is closest to the masked answer. For instance, in the examples shown in Fig. G, it can be observed that in both cases, GPT-4o chooses an option that is similar to the correct answer. The second is that there are cases where the correct answer is reached within the reasoning process but the final answer is incorrect. For example, in the example above in Fig. I, although the reasoning process mentions a predatory relationship, it is finally pulled towards a competitive relationship and answers “A". When we look up the meanings of “predatory relationship" and “competitive relationship" in a dictionary, we see that they are clearly different. Also, when we ask GPT-4o itself, it introduces them as different concepts. Therefore, this mistake is unique to UPD, and it shows the difficulty of refraining from answering. In the example below Fig. I, the reasoning stated the correct answer, “the magnitude of the magnetic force is greater in Pair 2. T", but GPT-4o chose “A" as a final answer. This also shows the difficulty of refraining from answering.
F.2 Qualitative Differences in Outputs Between Closed and Open Models
We compare some correct cases of GPT-4o, Gemini1.5Pro, LLaVA-NeXT-34B, and InternVL2-40B in Fig, J. Closed-source models often provide both the correct answer and an explanation like “None of the provided options are correct. The correct answer is …". In contrast, Open-source models typically only give the correct answer without providing “None of the ….". While both are considered correct in our evaluation, closed-source models offer a better response. The development of open-source models that can both provide the correct answer and respond with “None" is a crucial challenge for the future.
F.3 Other Failure Examples
Appendix G Full Results for Each Setting
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #10 | #11 | #12 | #13 | #14 | #15 | #16 | #17 | #18 | |
| Dual Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 2.0 | 0.0 | 0.0 | 0.0 | 4.3 | 8.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 3.4 | 0.0 | 0.0 | 0.0 |
| LLaVA-NeXT-13B | 8.2 | 0.0 | 20.0 | 35.1 | 34.8 | 12.0 | 31.6 | 23.8 | 2.2 | 21.7 | 0.0 | 42.2 | 3.2 | 20.0 | 3.4 | 0.0 | 37.5 | 11.4 |
| LLaVA-NeXT-34B | 57.1 | 29.5 | 58.0 | 68.1 | 60.9 | 20.0 | 81.6 | 66.7 | 59.1 | 45.7 | 38.7 | 66.7 | 21.0 | 62.9 | 10.3 | 20.0 | 56.2 | 25.0 |
| LLaVA-OV-0.5B | 2.0 | 2.3 | 38.0 | 29.8 | 41.3 | 0.0 | 76.3 | 52.4 | 21.5 | 8.7 | 9.7 | 2.2 | 8.1 | 31.4 | 0.0 | 6.7 | 53.1 | 2.3 |
| LLaVA-OV-7B | 2.0 | 0.0 | 2.0 | 27.7 | 0.0 | 0.0 | 0.0 | 0.0 | 2.2 | 0.0 | 9.7 | 0.0 | 0.0 | 5.7 | 0.0 | 0.0 | 0.0 | 4.5 |
| CogVLM-17B | 0.0 | 0.0 | 2.0 | 0.0 | 0.0 | 0.0 | 5.3 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.6 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| CogVLM2-19B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics2-8B | 4.1 | 0.0 | 4.0 | 4.3 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics3-8B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Phi3V | 0.0 | 0.0 | 2.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Phi3.5V | 2.0 | 0.0 | 2.0 | 2.1 | 0.0 | 0.0 | 13.2 | 4.8 | 3.2 | 0.0 | 3.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| InternVL2-2B | 2.0 | 0.0 | 20.0 | 10.6 | 4.3 | 0.0 | 0.0 | 31.0 | 0.0 | 0.0 | 9.7 | 2.2 | 0.0 | 11.4 | 3.4 | 0.0 | 21.9 | 9.1 |
| InternVL2-8B | 36.7 | 4.5 | 44.0 | 33.0 | 37.0 | 8.0 | 57.9 | 45.2 | 20.4 | 21.7 | 9.7 | 26.7 | 17.7 | 68.6 | 17.2 | 6.7 | 37.5 | 9.1 |
| InternVL2-40B | 51.0 | 9.1 | 60.0 | 71.3 | 47.8 | 16.0 | 78.9 | 54.8 | 31.2 | 10.9 | 41.9 | 35.6 | 40.3 | 68.6 | 44.8 | 6.7 | 53.1 | 20.5 |
| XgenMM | 10.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 2.6 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Qwen2-VL | 6.1 | 2.3 | 10.0 | 27.7 | 13.0 | 8.0 | 2.6 | 19.0 | 15.1 | 4.3 | 3.2 | 0.0 | 9.7 | 42.9 | 3.4 | 6.7 | 3.1 | 2.3 |
| Qwen2.5-VL | 65.3 | 29.5 | 22.0 | 60.6 | 39.1 | 12.0 | 28.9 | 2.4 | 46.2 | 17.4 | 45.2 | 22.2 | 30.6 | 28.6 | 6.9 | 13.3 | 12.5 | 13.6 |
| GeminiPro | 57.1 | 0.0 | 12.0 | 57.4 | 37.0 | 8.0 | 47.4 | 26.2 | 40.9 | 0.0 | 22.6 | 4.4 | 11.3 | 20.0 | 3.4 | 0.0 | 9.4 | 0.0 |
| Gemini1.5Pro | 71.4 | 4.5 | 60.0 | 72.3 | 52.2 | 28.0 | 89.5 | 28.6 | 79.6 | 30.4 | 71.0 | 31.1 | 9.7 | 82.9 | 3.4 | 0.0 | 18.8 | 31.8 |
| GPT4V | 85.7 | 2.3 | 52.0 | 48.9 | 63.0 | 16.0 | 92.1 | 50.0 | 93.5 | 41.3 | 77.4 | 66.7 | 11.3 | 88.6 | 10.3 | 6.7 | 21.9 | 38.6 |
| GPT4o-mini | 36.7 | 2.3 | 40.0 | 39.4 | 32.6 | 8.0 | 81.6 | 26.2 | 67.7 | 2.2 | 58.1 | 22.2 | 17.7 | 60.0 | 10.3 | 0.0 | 12.5 | 20.5 |
| GPT4o | 83.7 | 6.8 | 58.0 | 45.7 | 37.0 | 24.0 | 86.8 | 33.3 | 81.7 | 10.9 | 61.3 | 51.1 | 11.3 | 71.4 | 20.7 | 6.7 | 21.9 | 43.2 |
| UPD Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 2.0 | 0.0 | 0.0 | 0.0 | 4.3 | 16.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 3.4 | 0.0 | 0.0 | 0.0 |
| LLaVA-NeXT-13B | 8.2 | 0.0 | 22.0 | 36.2 | 39.1 | 44.0 | 31.6 | 23.8 | 2.2 | 28.3 | 0.0 | 55.6 | 3.2 | 20.0 | 10.3 | 20.0 | 78.1 | 22.7 |
| LLaVA-NeXT-34B | 65.3 | 31.8 | 66.0 | 74.5 | 71.7 | 88.0 | 94.7 | 85.7 | 59.1 | 56.5 | 38.7 | 73.3 | 25.8 | 62.9 | 27.6 | 20.0 | 87.5 | 31.8 |
| LLaVA-OV-0.5B | 2.0 | 2.3 | 50.0 | 33.0 | 45.7 | 4.0 | 78.9 | 54.8 | 21.5 | 19.6 | 9.7 | 2.2 | 16.1 | 31.4 | 3.4 | 6.7 | 59.4 | 4.5 |
| LLaVA-OV-7B | 2.0 | 0.0 | 2.0 | 30.9 | 0.0 | 0.0 | 0.0 | 0.0 | 2.2 | 2.2 | 9.7 | 0.0 | 0.0 | 5.7 | 0.0 | 0.0 | 0.0 | 6.8 |
| CogVLM-17B | 0.0 | 0.0 | 2.0 | 0.0 | 0.0 | 0.0 | 5.3 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.6 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| CogVLM2-19B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics2-8B | 4.1 | 0.0 | 4.0 | 4.3 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 2.3 |
| idefics3-8B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Phi3V | 0.0 | 0.0 | 2.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Phi3.5V | 2.0 | 0.0 | 2.0 | 2.1 | 0.0 | 0.0 | 13.2 | 4.8 | 3.2 | 0.0 | 3.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| InternVL2-2B | 2.0 | 0.0 | 20.0 | 10.6 | 4.3 | 0.0 | 0.0 | 31.0 | 0.0 | 0.0 | 9.7 | 2.2 | 1.6 | 11.4 | 3.4 | 0.0 | 21.9 | 11.4 |
| InternVL2-8B | 36.7 | 4.5 | 44.0 | 34.0 | 43.5 | 12.0 | 63.2 | 45.2 | 20.4 | 23.9 | 9.7 | 26.7 | 24.2 | 71.4 | 17.2 | 6.7 | 37.5 | 11.4 |
| InternVL2-40B | 51.0 | 13.6 | 60.0 | 74.5 | 52.2 | 24.0 | 78.9 | 54.8 | 31.2 | 10.9 | 41.9 | 35.6 | 45.2 | 71.4 | 44.8 | 13.3 | 53.1 | 27.3 |
| XgenMM | 10.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 2.6 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Qwen2-VL | 6.1 | 2.3 | 10.0 | 27.7 | 13.0 | 16.0 | 2.6 | 19.0 | 15.1 | 4.3 | 3.2 | 0.0 | 12.9 | 45.7 | 3.4 | 6.7 | 3.1 | 2.3 |
| Qwen2.5-VL | 65.3 | 31.8 | 24.0 | 61.7 | 43.5 | 24.0 | 28.9 | 2.4 | 46.2 | 17.4 | 45.2 | 26.7 | 32.3 | 28.6 | 6.9 | 13.3 | 12.5 | 18.2 |
| GeminiPro | 59.2 | 0.0 | 12.0 | 60.6 | 39.1 | 12.0 | 50.0 | 26.2 | 40.9 | 0.0 | 22.6 | 4.4 | 17.7 | 22.9 | 6.9 | 0.0 | 15.6 | 0.0 |
| Gemini1.5Pro | 87.8 | 4.5 | 64.0 | 73.4 | 56.5 | 56.0 | 89.5 | 31.0 | 86.0 | 34.8 | 71.0 | 33.3 | 17.7 | 85.7 | 6.9 | 13.3 | 18.8 | 34.1 |
| GPT4V | 91.8 | 4.5 | 56.0 | 89.4 | 63.0 | 76.0 | 94.7 | 50.0 | 93.5 | 45.7 | 77.4 | 68.9 | 17.7 | 91.4 | 34.5 | 46.7 | 40.6 | 43.2 |
| GPT4o-mini | 36.7 | 2.3 | 42.0 | 81.9 | 32.6 | 40.0 | 81.6 | 31.0 | 67.7 | 2.2 | 58.1 | 22.2 | 25.8 | 65.7 | 24.1 | 33.3 | 46.9 | 25.0 |
| GPT4o | 83.7 | 6.8 | 58.0 | 88.3 | 37.0 | 40.0 | 86.8 | 33.3 | 81.7 | 15.2 | 61.3 | 51.1 | 12.9 | 74.3 | 31.0 | 26.7 | 21.9 | 45.5 |
| Standard Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 95.9 | 59.1 | 92.0 | 84.0 | 87.0 | 52.0 | 100.0 | 85.7 | 97.8 | 69.6 | 90.3 | 44.4 | 50.0 | 62.9 | 13.8 | 40.0 | 87.5 | 18.2 |
| LLaVA-NeXT-13B | 93.9 | 59.1 | 86.0 | 86.2 | 80.4 | 40.0 | 100.0 | 85.7 | 97.8 | 82.6 | 93.5 | 55.6 | 53.2 | 71.4 | 6.9 | 46.7 | 56.2 | 22.7 |
| LLaVA-NeXT-34B | 79.6 | 54.5 | 72.0 | 80.9 | 78.3 | 20.0 | 86.8 | 66.7 | 97.8 | 73.9 | 100.0 | 84.4 | 64.5 | 82.9 | 17.2 | 40.0 | 65.6 | 50.0 |
| LLaVA-OV-0.5B | 93.9 | 4.5 | 66.0 | 59.6 | 80.4 | 8.0 | 94.7 | 81.0 | 96.8 | 52.2 | 87.1 | 37.8 | 33.9 | 62.9 | 0.0 | 33.3 | 75.0 | 6.8 |
| LLaVA-OV-7B | 100.0 | 45.5 | 96.0 | 88.3 | 97.8 | 68.0 | 100.0 | 90.5 | 98.9 | 91.3 | 100.0 | 84.4 | 69.4 | 94.3 | 69.0 | 86.7 | 96.9 | 43.2 |
| CogVLM-17B | 91.8 | 43.2 | 82.0 | 87.2 | 91.3 | 16.0 | 100.0 | 83.3 | 97.8 | 87.0 | 80.6 | 57.8 | 29.0 | 74.3 | 24.1 | 13.3 | 90.6 | 6.8 |
| CogVLM2-19B | 98.0 | 68.2 | 92.0 | 94.7 | 97.8 | 52.0 | 100.0 | 90.5 | 97.8 | 95.7 | 100.0 | 77.8 | 61.3 | 80.0 | 58.6 | 80.0 | 87.5 | 27.3 |
| idefics2-8B | 100.0 | 63.6 | 96.0 | 91.5 | 93.5 | 52.0 | 100.0 | 83.3 | 97.8 | 87.0 | 87.1 | 80.0 | 51.6 | 62.9 | 34.5 | 73.3 | 93.8 | 25.0 |
| idefics3-8B | 95.9 | 63.6 | 94.0 | 86.2 | 95.7 | 56.0 | 92.1 | 88.1 | 96.8 | 87.0 | 100.0 | 73.3 | 53.2 | 71.4 | 37.9 | 73.3 | 87.5 | 34.1 |
| Phi3V | 89.8 | 72.7 | 90.0 | 84.0 | 93.5 | 36.0 | 100.0 | 92.9 | 97.8 | 89.1 | 100.0 | 62.2 | 62.9 | 80.0 | 48.3 | 93.3 | 87.5 | 43.2 |
| Phi3.5V | 87.8 | 70.5 | 88.0 | 81.9 | 95.7 | 44.0 | 100.0 | 88.1 | 100.0 | 95.7 | 93.5 | 68.9 | 56.5 | 74.3 | 48.3 | 66.7 | 90.6 | 47.7 |
| InternVL2-2B | 98.0 | 38.6 | 88.0 | 78.7 | 80.4 | 40.0 | 97.4 | 88.1 | 100.0 | 84.8 | 93.5 | 35.6 | 58.1 | 88.6 | 41.4 | 86.7 | 93.8 | 27.3 |
| InternVL2-8B | 95.9 | 77.3 | 96.0 | 87.2 | 84.8 | 44.0 | 94.7 | 90.5 | 100.0 | 82.6 | 93.5 | 73.3 | 56.5 | 91.4 | 65.5 | 93.3 | 93.8 | 45.5 |
| InternVL2-40B | 95.9 | 72.7 | 96.0 | 95.7 | 95.7 | 56.0 | 100.0 | 88.1 | 97.8 | 89.1 | 93.5 | 95.6 | 71.0 | 91.4 | 65.5 | 73.3 | 93.8 | 59.1 |
| XgenMM | 91.8 | 81.8 | 96.0 | 92.6 | 97.8 | 52.0 | 100.0 | 90.5 | 96.8 | 91.3 | 93.5 | 62.2 | 69.4 | 77.1 | 34.5 | 80.0 | 90.6 | 27.3 |
| Qwen2-VL | 95.9 | 65.9 | 94.0 | 95.7 | 97.8 | 44.0 | 100.0 | 85.7 | 97.8 | 87.0 | 93.5 | 77.8 | 59.7 | 88.6 | 58.6 | 46.7 | 93.8 | 47.7 |
| Qwen2.5-VL | 95.9 | 70.5 | 94.0 | 96.8 | 93.5 | 56.0 | 100.0 | 88.1 | 97.8 | 91.3 | 100.0 | 73.3 | 64.5 | 97.1 | 58.6 | 53.3 | 93.8 | 70.5 |
| GeminiPro | 91.8 | 25.0 | 70.0 | 87.2 | 93.5 | 28.0 | 97.4 | 83.3 | 94.6 | 69.6 | 90.3 | 60.0 | 27.4 | 71.4 | 24.1 | 13.3 | 78.1 | 47.7 |
| Gemini1.5Pro | 77.6 | 54.5 | 78.0 | 93.6 | 87.0 | 44.0 | 97.4 | 83.3 | 91.4 | 87.0 | 90.3 | 91.1 | 38.7 | 94.3 | 34.5 | 33.3 | 62.5 | 77.3 |
| GPT4V | 91.8 | 45.5 | 82.0 | 56.4 | 97.8 | 28.0 | 97.4 | 88.1 | 97.8 | 93.5 | 100.0 | 93.3 | 40.3 | 94.3 | 31.0 | 20.0 | 56.2 | 72.7 |
| GPT4o-mini | 95.9 | 38.6 | 80.0 | 56.4 | 95.7 | 32.0 | 100.0 | 88.1 | 97.8 | 87.0 | 90.3 | 84.4 | 38.7 | 91.4 | 37.9 | 20.0 | 34.4 | 68.2 |
| GPT4o | 98.0 | 63.6 | 90.0 | 56.4 | 100.0 | 56.0 | 100.0 | 85.7 | 100.0 | 87.0 | 96.8 | 100.0 | 46.8 | 94.3 | 51.7 | 26.7 | 87.5 | 75.0 |
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #10 | #11 | #12 | #13 | #14 | #15 | #16 | #17 | #18 | |
| Dual Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 61.2 | 0.0 | 64.0 | 44.7 | 30.4 | 32.0 | 94.7 | 57.1 | 83.9 | 17.4 | 54.8 | 6.7 | 12.9 | 28.6 | 0.0 | 6.7 | 15.6 | 4.5 |
| LLaVA-NeXT-13B | 59.2 | 0.0 | 16.0 | 4.3 | 4.3 | 8.0 | 68.4 | 38.1 | 41.9 | 8.7 | 48.4 | 0.0 | 4.8 | 2.9 | 0.0 | 0.0 | 0.0 | 0.0 |
| LLaVA-NeXT-34B | 75.5 | 4.5 | 22.0 | 35.1 | 8.7 | 12.0 | 78.9 | 31.0 | 69.9 | 10.9 | 64.5 | 4.4 | 9.7 | 17.1 | 6.9 | 13.3 | 6.2 | 4.5 |
| LLaVA-OV-0.5B | 26.5 | 0.0 | 8.0 | 27.7 | 6.5 | 0.0 | 68.4 | 0.0 | 55.9 | 4.3 | 48.4 | 0.0 | 3.2 | 2.9 | 0.0 | 0.0 | 15.6 | 0.0 |
| LLaVA-OV-7B | 71.4 | 6.8 | 22.0 | 26.6 | 21.7 | 4.0 | 84.2 | 42.9 | 75.3 | 13.0 | 54.8 | 0.0 | 11.3 | 5.7 | 10.3 | 6.7 | 0.0 | 0.0 |
| CogVLM-17B | 79.6 | 0.0 | 48.0 | 50.0 | 52.2 | 0.0 | 86.8 | 61.9 | 65.6 | 30.4 | 51.6 | 26.7 | 4.8 | 25.7 | 6.9 | 0.0 | 37.5 | 0.0 |
| CogVLM2-19B | 77.6 | 0.0 | 62.0 | 70.2 | 37.0 | 16.0 | 86.8 | 57.1 | 88.2 | 41.3 | 64.5 | 26.7 | 12.9 | 25.7 | 10.3 | 6.7 | 25.0 | 6.8 |
| idefics2-8B | 83.7 | 0.0 | 42.0 | 17.0 | 28.3 | 8.0 | 71.1 | 38.1 | 76.3 | 17.4 | 48.4 | 4.4 | 6.5 | 5.7 | 3.4 | 6.7 | 15.6 | 4.5 |
| idefics3-8B | 87.8 | 9.1 | 62.0 | 28.7 | 23.9 | 16.0 | 71.1 | 42.9 | 72.0 | 13.0 | 51.6 | 4.4 | 9.7 | 14.3 | 6.9 | 13.3 | 6.2 | 0.0 |
| Phi3V | 61.2 | 34.1 | 12.0 | 26.6 | 6.5 | 4.0 | 78.9 | 26.2 | 68.8 | 21.7 | 54.8 | 4.4 | 4.8 | 11.4 | 3.4 | 0.0 | 3.1 | 4.5 |
| Phi3.5V | 65.3 | 2.3 | 16.0 | 20.2 | 6.5 | 16.0 | 71.1 | 14.3 | 48.4 | 10.9 | 41.9 | 20.0 | 8.1 | 8.6 | 0.0 | 0.0 | 3.1 | 2.3 |
| InternVL2-2B | 75.5 | 0.0 | 30.0 | 28.7 | 23.9 | 12.0 | 73.7 | 33.3 | 71.0 | 17.4 | 58.1 | 11.1 | 1.6 | 17.1 | 6.9 | 33.3 | 12.5 | 2.3 |
| InternVL2-8B | 85.7 | 27.3 | 74.0 | 63.8 | 69.6 | 36.0 | 94.7 | 47.6 | 93.5 | 56.5 | 71.0 | 22.2 | 17.7 | 45.7 | 17.2 | 20.0 | 71.9 | 18.2 |
| InternVL2-40B | 95.9 | 13.6 | 74.0 | 76.6 | 54.3 | 24.0 | 89.5 | 50.0 | 88.2 | 50.0 | 74.2 | 33.3 | 17.7 | 51.4 | 27.6 | 60.0 | 46.9 | 13.6 |
| XgenMM | 81.6 | 13.6 | 36.0 | 48.9 | 28.3 | 16.0 | 84.2 | 38.1 | 82.8 | 30.4 | 51.6 | 8.9 | 8.1 | 25.7 | 3.4 | 46.7 | 15.6 | 2.3 |
| Qwen2-VL | 89.8 | 4.5 | 46.0 | 36.2 | 52.2 | 12.0 | 86.8 | 50.0 | 81.7 | 26.1 | 58.1 | 20.0 | 9.7 | 20.0 | 3.4 | 6.7 | 0.0 | 2.3 |
| Qwen2.5-VL | 85.7 | 6.8 | 60.0 | 72.3 | 60.9 | 32.0 | 89.5 | 38.1 | 88.2 | 34.8 | 64.5 | 24.4 | 8.1 | 51.4 | 3.4 | 13.3 | 34.4 | 15.9 |
| GeminiPro | 87.8 | 2.3 | 36.0 | 70.2 | 26.1 | 8.0 | 78.9 | 40.5 | 75.3 | 26.1 | 64.5 | 22.2 | 11.3 | 34.3 | 6.9 | 0.0 | 18.8 | 2.3 |
| Gemini1.5Pro | 69.4 | 4.5 | 56.0 | 74.5 | 58.7 | 28.0 | 97.4 | 45.2 | 81.7 | 41.3 | 48.4 | 40.0 | 9.7 | 48.6 | 10.3 | 6.7 | 28.1 | 31.8 |
| GPT4V | 83.7 | 0.0 | 48.0 | 74.5 | 60.9 | 20.0 | 89.5 | 57.1 | 92.5 | 54.3 | 67.7 | 31.1 | 8.1 | 60.0 | 0.0 | 6.7 | 12.5 | 25.0 |
| GPT4o-mini | 81.6 | 0.0 | 62.0 | 69.1 | 50.0 | 24.0 | 97.4 | 45.2 | 89.2 | 43.5 | 64.5 | 31.1 | 8.1 | 54.3 | 0.0 | 13.3 | 15.6 | 27.3 |
| GPT4o | 93.9 | 31.8 | 74.0 | 48.9 | 65.2 | 32.0 | 94.7 | 66.7 | 97.8 | 50.0 | 67.7 | 48.9 | 9.7 | 68.6 | 10.3 | 6.7 | 68.8 | 36.4 |
| UPD Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 87.8 | 0.0 | 68.0 | 50.0 | 30.4 | 56.0 | 97.4 | 57.1 | 84.9 | 17.4 | 54.8 | 6.7 | 21.0 | 28.6 | 10.3 | 46.7 | 21.9 | 29.5 |
| LLaVA-NeXT-13B | 59.2 | 0.0 | 16.0 | 4.3 | 4.3 | 20.0 | 68.4 | 38.1 | 41.9 | 8.7 | 48.4 | 0.0 | 4.8 | 2.9 | 0.0 | 0.0 | 0.0 | 0.0 |
| LLaVA-NeXT-34B | 75.5 | 4.5 | 22.0 | 35.1 | 8.7 | 16.0 | 78.9 | 31.0 | 69.9 | 10.9 | 64.5 | 6.7 | 9.7 | 17.1 | 6.9 | 13.3 | 6.2 | 4.5 |
| LLaVA-OV-0.5B | 26.5 | 0.0 | 8.0 | 30.9 | 6.5 | 0.0 | 68.4 | 0.0 | 55.9 | 4.3 | 48.4 | 0.0 | 8.1 | 2.9 | 0.0 | 0.0 | 15.6 | 0.0 |
| LLaVA-OV-7B | 71.4 | 6.8 | 22.0 | 27.7 | 21.7 | 8.0 | 84.2 | 42.9 | 75.3 | 13.0 | 54.8 | 0.0 | 11.3 | 5.7 | 10.3 | 6.7 | 0.0 | 0.0 |
| CogVLM-17B | 81.6 | 0.0 | 50.0 | 50.0 | 52.2 | 0.0 | 86.8 | 64.3 | 65.6 | 32.6 | 58.1 | 48.9 | 6.5 | 25.7 | 13.8 | 0.0 | 37.5 | 15.9 |
| CogVLM2-19B | 77.6 | 0.0 | 64.0 | 71.3 | 39.1 | 20.0 | 86.8 | 61.9 | 88.2 | 41.3 | 64.5 | 31.1 | 17.7 | 25.7 | 10.3 | 6.7 | 25.0 | 9.1 |
| idefics2-8B | 91.8 | 0.0 | 44.0 | 17.0 | 30.4 | 8.0 | 78.9 | 40.5 | 80.6 | 19.6 | 51.6 | 6.7 | 9.7 | 5.7 | 3.4 | 6.7 | 15.6 | 6.8 |
| idefics3-8B | 91.8 | 13.6 | 64.0 | 31.9 | 23.9 | 32.0 | 78.9 | 42.9 | 74.2 | 15.2 | 54.8 | 4.4 | 14.5 | 14.3 | 10.3 | 13.3 | 6.2 | 2.3 |
| Phi3V | 69.4 | 40.9 | 12.0 | 27.7 | 8.7 | 40.0 | 78.9 | 28.6 | 69.9 | 21.7 | 58.1 | 4.4 | 8.1 | 11.4 | 6.9 | 0.0 | 3.1 | 20.5 |
| Phi3.5V | 67.3 | 2.3 | 16.0 | 23.4 | 6.5 | 28.0 | 71.1 | 16.7 | 48.4 | 10.9 | 41.9 | 26.7 | 11.3 | 8.6 | 0.0 | 0.0 | 3.1 | 2.3 |
| InternVL2-2B | 75.5 | 0.0 | 30.0 | 30.9 | 30.4 | 28.0 | 73.7 | 33.3 | 71.0 | 17.4 | 58.1 | 15.6 | 9.7 | 17.1 | 10.3 | 33.3 | 15.6 | 4.5 |
| InternVL2-8B | 98.0 | 29.5 | 74.0 | 64.9 | 71.7 | 52.0 | 94.7 | 47.6 | 94.6 | 60.9 | 74.2 | 26.7 | 21.0 | 45.7 | 17.2 | 20.0 | 75.0 | 27.3 |
| InternVL2-40B | 98.0 | 13.6 | 74.0 | 76.6 | 54.3 | 32.0 | 89.5 | 50.0 | 88.2 | 52.2 | 74.2 | 37.8 | 17.7 | 51.4 | 27.6 | 60.0 | 46.9 | 13.6 |
| XgenMM | 89.8 | 15.9 | 38.0 | 51.1 | 28.3 | 20.0 | 86.8 | 38.1 | 82.8 | 32.6 | 61.3 | 15.6 | 11.3 | 25.7 | 10.3 | 46.7 | 18.8 | 6.8 |
| Qwen2-VL | 89.8 | 9.1 | 50.0 | 37.2 | 54.3 | 24.0 | 86.8 | 50.0 | 81.7 | 26.1 | 58.1 | 22.2 | 11.3 | 20.0 | 3.4 | 6.7 | 0.0 | 4.5 |
| Qwen2.5-VL | 93.9 | 9.1 | 60.0 | 72.3 | 63.0 | 52.0 | 89.5 | 38.1 | 88.2 | 34.8 | 64.5 | 40.0 | 9.7 | 51.4 | 3.4 | 20.0 | 34.4 | 31.8 |
| GeminiPro | 87.8 | 13.6 | 52.0 | 71.3 | 32.6 | 20.0 | 78.9 | 50.0 | 77.4 | 34.8 | 64.5 | 33.3 | 32.3 | 34.3 | 31.0 | 6.7 | 31.2 | 2.3 |
| Gemini1.5Pro | 93.9 | 6.8 | 76.0 | 78.7 | 67.4 | 64.0 | 97.4 | 54.8 | 94.6 | 54.3 | 71.0 | 44.4 | 33.9 | 51.4 | 20.7 | 40.0 | 28.1 | 31.8 |
| GPT4V | 100.0 | 0.0 | 64.0 | 89.4 | 67.4 | 88.0 | 92.1 | 69.0 | 93.5 | 60.9 | 67.7 | 35.6 | 14.5 | 62.9 | 20.7 | 53.3 | 18.8 | 29.5 |
| GPT4o-mini | 93.9 | 2.3 | 78.0 | 88.3 | 56.5 | 48.0 | 97.4 | 54.8 | 92.5 | 50.0 | 67.7 | 31.1 | 22.6 | 54.3 | 6.9 | 20.0 | 28.1 | 27.3 |
| GPT4o | 95.9 | 31.8 | 82.0 | 91.5 | 71.7 | 56.0 | 97.4 | 78.6 | 98.9 | 54.3 | 74.2 | 53.3 | 14.5 | 68.6 | 10.3 | 6.7 | 68.8 | 38.6 |
| Standard Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 67.3 | 61.4 | 88.0 | 79.8 | 89.1 | 56.0 | 97.4 | 88.1 | 97.8 | 76.1 | 90.3 | 51.1 | 46.8 | 60.0 | 6.9 | 33.3 | 87.5 | 13.6 |
| LLaVA-NeXT-13B | 87.8 | 61.4 | 86.0 | 84.0 | 89.1 | 60.0 | 100.0 | 88.1 | 97.8 | 87.0 | 93.5 | 60.0 | 50.0 | 68.6 | 27.6 | 53.3 | 93.8 | 25.0 |
| LLaVA-NeXT-34B | 91.8 | 75.0 | 90.0 | 88.3 | 93.5 | 68.0 | 97.4 | 90.5 | 98.9 | 93.5 | 96.8 | 84.4 | 62.9 | 82.9 | 48.3 | 46.7 | 96.9 | 59.1 |
| LLaVA-OV-0.5B | 93.9 | 0.0 | 84.0 | 71.3 | 89.1 | 12.0 | 100.0 | 90.5 | 100.0 | 78.3 | 90.3 | 40.0 | 46.8 | 68.6 | 3.4 | 33.3 | 84.4 | 6.8 |
| LLaVA-OV-7B | 100.0 | 43.2 | 96.0 | 90.4 | 97.8 | 60.0 | 100.0 | 90.5 | 98.9 | 91.3 | 100.0 | 82.2 | 67.7 | 94.3 | 65.5 | 86.7 | 96.9 | 54.5 |
| CogVLM-17B | 93.9 | 56.8 | 84.0 | 87.2 | 89.1 | 36.0 | 100.0 | 88.1 | 96.8 | 87.0 | 80.6 | 53.3 | 17.7 | 71.4 | 24.1 | 13.3 | 90.6 | 4.5 |
| CogVLM2-19B | 98.0 | 70.5 | 92.0 | 92.6 | 97.8 | 52.0 | 100.0 | 88.1 | 97.8 | 95.7 | 100.0 | 80.0 | 58.1 | 80.0 | 62.1 | 80.0 | 87.5 | 27.3 |
| idefics2-8B | 91.8 | 56.8 | 92.0 | 90.4 | 91.3 | 52.0 | 84.2 | 85.7 | 92.5 | 84.8 | 83.9 | 77.8 | 48.4 | 57.1 | 13.8 | 66.7 | 87.5 | 27.3 |
| idefics3-8B | 93.9 | 65.9 | 94.0 | 87.2 | 100.0 | 52.0 | 89.5 | 88.1 | 95.7 | 84.8 | 96.8 | 73.3 | 53.2 | 74.3 | 41.4 | 80.0 | 93.8 | 25.0 |
| Phi3V | 83.7 | 70.5 | 90.0 | 84.0 | 91.3 | 32.0 | 100.0 | 90.5 | 96.8 | 84.8 | 96.8 | 60.0 | 59.7 | 80.0 | 44.8 | 93.3 | 90.6 | 25.0 |
| Phi3.5V | 87.8 | 70.5 | 92.0 | 77.7 | 95.7 | 52.0 | 100.0 | 88.1 | 100.0 | 95.7 | 93.5 | 62.2 | 56.5 | 74.3 | 41.4 | 73.3 | 90.6 | 50.0 |
| InternVL2-2B | 98.0 | 40.9 | 94.0 | 85.1 | 82.6 | 44.0 | 100.0 | 88.1 | 100.0 | 84.8 | 93.5 | 42.2 | 62.9 | 77.1 | 72.4 | 86.7 | 90.6 | 29.5 |
| InternVL2-8B | 85.7 | 61.4 | 94.0 | 93.6 | 95.7 | 64.0 | 100.0 | 90.5 | 97.8 | 87.0 | 93.5 | 73.3 | 79.0 | 88.6 | 72.4 | 80.0 | 90.6 | 56.8 |
| InternVL2-40B | 98.0 | 81.8 | 98.0 | 97.9 | 97.8 | 64.0 | 100.0 | 88.1 | 98.9 | 91.3 | 96.8 | 86.7 | 83.9 | 94.3 | 72.4 | 86.7 | 96.9 | 63.6 |
| XgenMM | 85.7 | 86.4 | 92.0 | 90.4 | 97.8 | 48.0 | 97.4 | 90.5 | 96.8 | 89.1 | 87.1 | 60.0 | 69.4 | 80.0 | 44.8 | 73.3 | 90.6 | 29.5 |
| Qwen2-VL | 95.9 | 70.5 | 94.0 | 94.7 | 95.7 | 60.0 | 100.0 | 88.1 | 98.9 | 95.7 | 93.5 | 77.8 | 66.1 | 85.7 | 41.4 | 53.3 | 96.9 | 47.7 |
| Qwen2.5-VL | 89.8 | 84.1 | 98.0 | 96.8 | 97.8 | 64.0 | 100.0 | 90.5 | 100.0 | 93.5 | 100.0 | 66.7 | 64.5 | 94.3 | 58.6 | 53.3 | 96.9 | 59.1 |
| GeminiPro | 95.9 | 36.4 | 64.0 | 89.4 | 84.8 | 44.0 | 97.4 | 78.6 | 94.6 | 73.9 | 87.1 | 60.0 | 32.3 | 74.3 | 24.1 | 20.0 | 87.5 | 52.3 |
| Gemini1.5Pro | 71.4 | 50.0 | 68.0 | 90.4 | 87.0 | 48.0 | 97.4 | 81.0 | 86.0 | 78.3 | 71.0 | 88.9 | 32.3 | 94.3 | 41.4 | 20.0 | 68.8 | 75.0 |
| GPT4V | 83.7 | 36.4 | 70.0 | 81.9 | 93.5 | 32.0 | 97.4 | 81.0 | 97.8 | 84.8 | 100.0 | 82.2 | 24.2 | 94.3 | 31.0 | 26.7 | 75.0 | 63.6 |
| GPT4o-mini | 87.8 | 52.3 | 72.0 | 77.7 | 91.3 | 48.0 | 100.0 | 85.7 | 94.6 | 84.8 | 87.1 | 75.6 | 27.4 | 94.3 | 24.1 | 20.0 | 68.8 | 63.6 |
| GPT4o | 95.9 | 65.9 | 84.0 | 57.4 | 93.5 | 60.0 | 97.4 | 81.0 | 98.9 | 89.1 | 93.5 | 95.6 | 43.5 | 97.1 | 55.2 | 33.3 | 90.6 | 84.1 |
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #10 | #11 | #12 | #13 | #14 | #15 | #16 | #17 | #18 | |
| Dual Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 59.2 | 0.0 | 46.0 | 33.0 | 58.7 | 36.0 | 89.5 | 47.6 | 72.0 | 17.4 | 54.8 | 11.1 | 9.7 | 28.6 | 3.4 | 13.3 | 43.8 | 2.3 |
| LLaVA-NeXT-13B | 63.3 | 0.0 | 50.0 | 37.2 | 52.2 | 20.0 | 81.6 | 57.1 | 63.4 | 34.8 | 48.4 | 28.9 | 19.4 | 28.6 | 0.0 | 13.3 | 25.0 | 9.1 |
| LLaVA-NeXT-34B | 51.0 | 34.1 | 78.0 | 57.4 | 65.2 | 16.0 | 76.3 | 73.8 | 80.6 | 58.7 | 64.5 | 62.2 | 29.0 | 62.9 | 10.3 | 33.3 | 75.0 | 6.8 |
| LLaVA-OV-0.5B | 0.0 | 0.0 | 0.0 | 0.0 | 2.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| LLaVA-OV-7B | 71.4 | 0.0 | 6.0 | 39.4 | 19.6 | 4.0 | 55.3 | 23.8 | 66.7 | 6.5 | 48.4 | 0.0 | 4.8 | 5.7 | 6.9 | 6.7 | 6.2 | 13.6 |
| CogVLM-17B | 0.0 | 0.0 | 12.0 | 7.4 | 8.7 | 0.0 | 0.0 | 0.0 | 4.3 | 6.5 | 6.5 | 6.7 | 0.0 | 5.7 | 0.0 | 0.0 | 0.0 | 0.0 |
| CogVLM2-19B | 65.3 | 0.0 | 52.0 | 64.9 | 50.0 | 20.0 | 86.8 | 57.1 | 84.9 | 43.5 | 67.7 | 20.0 | 14.5 | 31.4 | 3.4 | 0.0 | 28.1 | 4.5 |
| idefics2-8B | 69.4 | 0.0 | 30.0 | 13.8 | 28.3 | 12.0 | 71.1 | 40.5 | 79.6 | 15.2 | 48.4 | 0.0 | 3.2 | 2.9 | 0.0 | 6.7 | 0.0 | 4.5 |
| idefics3-8B | 91.8 | 2.3 | 42.0 | 24.5 | 21.7 | 16.0 | 57.9 | 31.0 | 65.6 | 8.7 | 58.1 | 11.1 | 6.5 | 5.7 | 3.4 | 13.3 | 6.2 | 2.3 |
| Phi3V | 65.3 | 38.6 | 34.0 | 39.4 | 30.4 | 16.0 | 86.8 | 42.9 | 81.7 | 34.8 | 67.7 | 6.7 | 8.1 | 31.4 | 6.9 | 20.0 | 18.8 | 6.8 |
| Phi3.5V | 61.2 | 0.0 | 26.0 | 37.2 | 15.2 | 24.0 | 84.2 | 31.0 | 51.6 | 17.4 | 45.2 | 8.9 | 8.1 | 14.3 | 6.9 | 0.0 | 9.4 | 4.5 |
| InternVL2-2B | 53.1 | 0.0 | 14.0 | 13.8 | 15.2 | 12.0 | 57.9 | 4.8 | 34.4 | 10.9 | 22.6 | 6.7 | 4.8 | 14.3 | 6.9 | 13.3 | 3.1 | 6.8 |
| InternVL2-8B | 81.6 | 4.5 | 30.0 | 29.8 | 47.8 | 20.0 | 68.4 | 28.6 | 75.3 | 6.5 | 54.8 | 4.4 | 4.8 | 11.4 | 27.6 | 13.3 | 50.0 | 9.1 |
| InternVL2-40B | 81.6 | 38.6 | 86.0 | 88.3 | 84.8 | 12.0 | 89.5 | 66.7 | 96.8 | 58.7 | 87.1 | 46.7 | 30.6 | 77.1 | 41.4 | 46.7 | 87.5 | 27.3 |
| XgenMM | 79.6 | 11.4 | 20.0 | 18.1 | 34.8 | 20.0 | 84.2 | 33.3 | 78.5 | 15.2 | 54.8 | 2.2 | 4.8 | 17.1 | 3.4 | 46.7 | 15.6 | 2.3 |
| Qwen2-VL | 89.8 | 2.3 | 52.0 | 71.3 | 67.4 | 28.0 | 92.1 | 57.1 | 87.1 | 28.3 | 58.1 | 24.4 | 9.7 | 42.9 | 13.8 | 13.3 | 18.8 | 11.4 |
| Qwen2.5-VL | 85.7 | 50.0 | 80.0 | 79.8 | 58.7 | 16.0 | 81.6 | 45.2 | 93.5 | 47.8 | 58.1 | 40.0 | 38.7 | 62.9 | 10.3 | 33.3 | 37.5 | 20.5 |
| GeminiPro | 79.6 | 0.0 | 46.0 | 74.5 | 52.2 | 16.0 | 92.1 | 47.6 | 71.0 | 28.3 | 58.1 | 13.3 | 9.7 | 40.0 | 13.8 | 6.7 | 25.0 | 2.3 |
| Gemini1.5Pro | 49.0 | 15.9 | 58.0 | 74.5 | 76.1 | 32.0 | 94.7 | 59.5 | 74.2 | 65.2 | 41.9 | 44.4 | 14.5 | 80.0 | 6.9 | 6.7 | 18.8 | 38.6 |
| GPT4V | 71.4 | 20.5 | 54.0 | 60.6 | 73.9 | 24.0 | 92.1 | 73.8 | 92.5 | 65.2 | 74.2 | 46.7 | 12.9 | 77.1 | 20.7 | 13.3 | 34.4 | 34.1 |
| GPT4o-mini | 77.6 | 2.3 | 56.0 | 42.6 | 52.2 | 12.0 | 86.8 | 42.9 | 78.5 | 41.3 | 61.3 | 40.0 | 6.5 | 77.1 | 17.2 | 13.3 | 15.6 | 29.5 |
| GPT4o | 93.9 | 20.5 | 74.0 | 50.0 | 71.7 | 20.0 | 89.5 | 64.3 | 97.8 | 52.2 | 61.3 | 68.9 | 9.7 | 85.7 | 27.6 | 6.7 | 59.4 | 43.2 |
| UPD Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 91.8 | 0.0 | 46.0 | 36.2 | 58.7 | 84.0 | 94.7 | 47.6 | 73.1 | 17.4 | 58.1 | 11.1 | 17.7 | 28.6 | 20.7 | 60.0 | 53.1 | 22.7 |
| LLaVA-NeXT-13B | 87.8 | 0.0 | 54.0 | 42.6 | 58.7 | 80.0 | 92.1 | 64.3 | 66.7 | 41.3 | 58.1 | 40.0 | 25.8 | 28.6 | 10.3 | 26.7 | 71.9 | 15.9 |
| LLaVA-NeXT-34B | 98.0 | 59.1 | 86.0 | 87.2 | 73.9 | 96.0 | 97.4 | 90.5 | 90.3 | 67.4 | 93.5 | 73.3 | 40.3 | 71.4 | 48.3 | 80.0 | 78.1 | 59.1 |
| LLaVA-OV-0.5B | 0.0 | 0.0 | 0.0 | 0.0 | 4.3 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| LLaVA-OV-7B | 71.4 | 0.0 | 6.0 | 41.5 | 19.6 | 12.0 | 55.3 | 23.8 | 66.7 | 6.5 | 51.6 | 0.0 | 4.8 | 5.7 | 6.9 | 6.7 | 6.2 | 29.5 |
| CogVLM-17B | 100.0 | 93.2 | 94.0 | 93.6 | 93.5 | 96.0 | 100.0 | 100.0 | 92.5 | 95.7 | 96.8 | 95.6 | 75.8 | 91.4 | 93.1 | 40.0 | 93.8 | 90.9 |
| CogVLM2-19B | 91.8 | 0.0 | 58.0 | 67.0 | 54.3 | 28.0 | 86.8 | 66.7 | 90.3 | 43.5 | 67.7 | 24.4 | 17.7 | 31.4 | 3.4 | 0.0 | 28.1 | 6.8 |
| idefics2-8B | 95.9 | 0.0 | 30.0 | 13.8 | 28.3 | 12.0 | 73.7 | 40.5 | 80.6 | 15.2 | 51.6 | 0.0 | 4.8 | 2.9 | 0.0 | 6.7 | 3.1 | 4.5 |
| idefics3-8B | 91.8 | 2.3 | 42.0 | 24.5 | 21.7 | 20.0 | 57.9 | 31.0 | 67.7 | 10.9 | 58.1 | 11.1 | 8.1 | 5.7 | 10.3 | 13.3 | 6.2 | 4.5 |
| Phi3V | 81.6 | 47.7 | 34.0 | 41.5 | 34.8 | 56.0 | 86.8 | 50.0 | 82.8 | 34.8 | 67.7 | 13.3 | 12.9 | 31.4 | 37.9 | 26.7 | 18.8 | 18.2 |
| Phi3.5V | 71.4 | 0.0 | 26.0 | 41.5 | 15.2 | 48.0 | 84.2 | 35.7 | 51.6 | 17.4 | 45.2 | 8.9 | 12.9 | 17.1 | 17.2 | 0.0 | 12.5 | 15.9 |
| InternVL2-2B | 53.1 | 0.0 | 14.0 | 13.8 | 15.2 | 16.0 | 57.9 | 4.8 | 34.4 | 10.9 | 22.6 | 6.7 | 4.8 | 14.3 | 13.8 | 13.3 | 3.1 | 6.8 |
| InternVLM2-8B | 85.7 | 6.8 | 30.0 | 29.8 | 52.2 | 28.0 | 68.4 | 28.6 | 75.3 | 6.5 | 54.8 | 4.4 | 4.8 | 11.4 | 27.6 | 13.3 | 50.0 | 15.9 |
| InternVL2-40B | 100.0 | 45.5 | 92.0 | 90.4 | 95.7 | 92.0 | 94.7 | 81.0 | 98.9 | 76.1 | 93.5 | 57.8 | 35.5 | 80.0 | 44.8 | 46.7 | 100.0 | 43.2 |
| XgenMM | 83.7 | 11.4 | 20.0 | 18.1 | 34.8 | 24.0 | 84.2 | 33.3 | 78.5 | 17.4 | 54.8 | 4.4 | 4.8 | 20.0 | 10.3 | 46.7 | 18.8 | 6.8 |
| Qwen2-VL | 95.9 | 4.5 | 52.0 | 74.5 | 67.4 | 52.0 | 92.1 | 59.5 | 87.1 | 28.3 | 58.1 | 28.9 | 12.9 | 48.6 | 13.8 | 20.0 | 18.8 | 22.7 |
| Qwen2.5-VL | 98.0 | 56.8 | 86.0 | 84.0 | 60.9 | 40.0 | 86.8 | 47.6 | 94.6 | 50.0 | 58.1 | 48.9 | 41.9 | 62.9 | 10.3 | 33.3 | 37.5 | 36.4 |
| GeminiPro | 83.7 | 4.5 | 64.0 | 84.0 | 58.7 | 68.0 | 97.4 | 54.8 | 76.3 | 43.5 | 67.7 | 28.9 | 48.4 | 42.9 | 24.1 | 20.0 | 40.6 | 2.3 |
| Gemini1.5Pro | 98.0 | 25.0 | 86.0 | 85.1 | 91.3 | 80.0 | 94.7 | 66.7 | 95.7 | 82.6 | 77.4 | 51.1 | 66.1 | 82.9 | 27.6 | 80.0 | 31.2 | 54.5 |
| GPT4V | 100.0 | 36.4 | 90.0 | 96.8 | 93.5 | 100.0 | 97.4 | 97.6 | 95.7 | 87.0 | 80.6 | 88.9 | 74.2 | 80.0 | 62.1 | 73.3 | 40.6 | 47.7 |
| GPT4o-mini | 93.9 | 18.2 | 82.0 | 91.5 | 58.7 | 84.0 | 97.4 | 57.1 | 89.2 | 56.5 | 71.0 | 48.9 | 37.1 | 80.0 | 51.7 | 53.3 | 75.0 | 38.6 |
| GPT4o | 95.9 | 20.5 | 82.0 | 94.7 | 73.9 | 80.0 | 97.4 | 81.0 | 100.0 | 60.9 | 71.0 | 80.0 | 19.4 | 85.7 | 41.4 | 46.7 | 68.8 | 47.7 |
| Standard Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 63.3 | 59.1 | 90.0 | 81.9 | 89.1 | 44.0 | 94.7 | 85.7 | 96.8 | 69.6 | 87.1 | 48.9 | 45.2 | 62.9 | 6.9 | 26.7 | 81.2 | 13.6 |
| LLaVA-NeXT-13B | 71.4 | 56.8 | 80.0 | 81.9 | 76.1 | 28.0 | 89.5 | 83.3 | 94.6 | 82.6 | 80.6 | 46.7 | 46.8 | 68.6 | 10.3 | 40.0 | 46.9 | 18.2 |
| LLaVA-NeXT-34B | 53.1 | 50.0 | 84.0 | 64.9 | 87.0 | 20.0 | 78.9 | 83.3 | 89.2 | 80.4 | 71.0 | 77.8 | 58.1 | 77.1 | 24.1 | 46.7 | 90.6 | 22.7 |
| LLaVA-OV-0.5B | 87.8 | 2.3 | 76.0 | 67.0 | 76.1 | 8.0 | 97.4 | 83.3 | 97.8 | 54.3 | 90.3 | 35.6 | 19.4 | 60.0 | 0.0 | 33.3 | 62.5 | 4.5 |
| LLaVA-OV-7B | 100.0 | 40.9 | 96.0 | 89.4 | 97.8 | 56.0 | 100.0 | 90.5 | 98.9 | 91.3 | 96.8 | 82.2 | 66.1 | 91.4 | 69.0 | 86.7 | 96.9 | 40.9 |
| CogVLM-17B | 0.0 | 0.0 | 14.0 | 8.5 | 13.0 | 0.0 | 0.0 | 0.0 | 9.7 | 6.5 | 9.7 | 6.7 | 0.0 | 14.3 | 0.0 | 0.0 | 0.0 | 0.0 |
| CogVLM2-19B | 73.5 | 65.9 | 88.0 | 90.4 | 95.7 | 52.0 | 100.0 | 85.7 | 92.5 | 93.5 | 100.0 | 84.4 | 59.7 | 80.0 | 55.2 | 80.0 | 87.5 | 25.0 |
| idefics2-8B | 71.4 | 63.6 | 92.0 | 91.5 | 93.5 | 48.0 | 97.4 | 88.1 | 96.8 | 87.0 | 83.9 | 80.0 | 48.4 | 62.9 | 37.9 | 60.0 | 84.4 | 22.7 |
| idefics3-8B | 98.0 | 68.2 | 98.0 | 87.2 | 100.0 | 56.0 | 100.0 | 90.5 | 95.7 | 87.0 | 100.0 | 73.3 | 54.8 | 71.4 | 37.9 | 80.0 | 93.8 | 31.8 |
| Phi3V | 77.6 | 68.2 | 90.0 | 84.0 | 91.3 | 24.0 | 100.0 | 88.1 | 96.8 | 87.0 | 100.0 | 55.6 | 59.7 | 80.0 | 37.9 | 86.7 | 87.5 | 34.1 |
| Phi3.5V | 81.6 | 72.7 | 88.0 | 77.7 | 93.5 | 40.0 | 100.0 | 88.1 | 98.9 | 93.5 | 93.5 | 68.9 | 51.6 | 71.4 | 44.8 | 73.3 | 90.6 | 36.4 |
| InternVL2-2B | 98.0 | 43.2 | 94.0 | 81.9 | 91.3 | 56.0 | 100.0 | 88.1 | 100.0 | 84.8 | 93.5 | 46.7 | 61.3 | 77.1 | 37.9 | 86.7 | 93.8 | 34.1 |
| InternVLM2-8B | 93.9 | 72.7 | 96.0 | 94.7 | 93.5 | 64.0 | 100.0 | 90.5 | 98.9 | 89.1 | 90.3 | 71.1 | 75.8 | 97.1 | 72.4 | 86.7 | 93.8 | 45.5 |
| InternVL2-40B | 81.6 | 75.0 | 92.0 | 95.7 | 89.1 | 12.0 | 94.7 | 83.3 | 96.8 | 76.1 | 93.5 | 80.0 | 75.8 | 94.3 | 75.9 | 86.7 | 87.5 | 63.6 |
| XgenMM | 89.8 | 75.0 | 94.0 | 92.6 | 97.8 | 48.0 | 100.0 | 90.5 | 96.8 | 89.1 | 96.8 | 60.0 | 69.4 | 74.3 | 31.0 | 73.3 | 90.6 | 27.3 |
| Qwen2-VL | 91.8 | 68.2 | 94.0 | 93.6 | 100.0 | 64.0 | 100.0 | 88.1 | 98.9 | 91.3 | 93.5 | 82.2 | 66.1 | 82.9 | 58.6 | 53.3 | 96.9 | 36.4 |
| Qwen2.5-VL | 85.7 | 70.5 | 86.0 | 93.6 | 93.5 | 44.0 | 94.7 | 85.7 | 96.8 | 91.3 | 100.0 | 75.6 | 64.5 | 97.1 | 62.1 | 53.3 | 93.8 | 52.3 |
| GeminiPro | 95.9 | 34.1 | 62.0 | 83.0 | 89.1 | 28.0 | 94.7 | 78.6 | 90.3 | 67.4 | 74.2 | 53.3 | 17.7 | 65.7 | 20.7 | 20.0 | 78.1 | 50.0 |
| Gemini1.5Pro | 49.0 | 38.6 | 64.0 | 87.2 | 80.4 | 44.0 | 97.4 | 83.3 | 76.3 | 76.1 | 58.1 | 88.9 | 27.4 | 91.4 | 24.1 | 13.3 | 50.0 | 59.1 |
| GPT4V | 71.4 | 40.9 | 58.0 | 62.8 | 78.3 | 24.0 | 94.7 | 73.8 | 95.7 | 76.1 | 93.5 | 55.6 | 19.4 | 94.3 | 34.5 | 13.3 | 75.0 | 59.1 |
| GPT4o-mini | 81.6 | 45.5 | 64.0 | 48.9 | 89.1 | 24.0 | 86.8 | 81.0 | 87.1 | 67.4 | 77.4 | 75.6 | 17.7 | 94.3 | 34.5 | 20.0 | 25.0 | 59.1 |
| GPT4o | 95.9 | 61.4 | 84.0 | 55.3 | 97.8 | 36.0 | 92.1 | 81.0 | 97.8 | 80.4 | 90.3 | 88.9 | 32.3 | 97.1 | 55.2 | 26.7 | 78.1 | 81.8 |
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #10 | #11 | #12 | #13 | #14 | #15 | #16 | #17 | #18 | |
| Dual Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 1.9 | 12.8 | 1.9 | 1.0 | 25.4 | 0.0 | 2.4 | 2.0 | 1.0 | 0.0 | 3.0 | 9.3 | 3.9 | 10.3 | 2.4 | 10.0 | 23.8 | 0.0 |
| LLaVA-NeXT-13B | 24.5 | 17.9 | 35.2 | 41.2 | 41.3 | 20.0 | 43.9 | 36.7 | 9.2 | 21.6 | 3.0 | 39.5 | 13.0 | 46.2 | 14.3 | 5.0 | 45.2 | 18.6 |
| LLaVA-NeXT-34B | 50.9 | 46.2 | 53.7 | 61.9 | 41.3 | 8.6 | 80.5 | 55.1 | 49.0 | 62.7 | 42.4 | 74.4 | 27.3 | 56.4 | 26.2 | 20.0 | 59.5 | 39.5 |
| LLaVA-OV-0.5B | 7.5 | 0.0 | 22.2 | 22.7 | 34.9 | 2.9 | 29.3 | 24.5 | 26.5 | 13.7 | 12.1 | 7.0 | 10.4 | 28.2 | 2.4 | 15.0 | 35.7 | 2.3 |
| LLaVA-OV-7B | 0.0 | 2.6 | 1.9 | 27.8 | 1.6 | 2.9 | 0.0 | 4.1 | 2.0 | 3.9 | 3.0 | 2.3 | 0.0 | 10.3 | 4.8 | 0.0 | 9.5 | 4.7 |
| CogVLM-17B | 0.0 | 5.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.3 | 2.6 | 0.0 | 0.0 | 2.4 | 0.0 |
| CogVLM2-19B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics2-8B | 0.0 | 0.0 | 1.9 | 4.1 | 0.0 | 2.9 | 0.0 | 4.1 | 1.0 | 0.0 | 3.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics3-8B | 0.0 | 2.6 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 4.8 | 0.0 | 0.0 | 0.0 |
| Phi3V | 0.0 | 0.0 | 1.9 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Phi3.5V | 0.0 | 0.0 | 1.9 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.0 | 0.0 | 3.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| InternVL2-2B | 1.9 | 10.3 | 37.0 | 15.5 | 27.0 | 0.0 | 12.2 | 28.6 | 3.1 | 5.9 | 9.1 | 4.7 | 6.5 | 43.6 | 4.8 | 30.0 | 38.1 | 2.3 |
| InternVL2-8B | 20.8 | 33.3 | 42.6 | 41.2 | 36.5 | 14.3 | 36.6 | 28.6 | 31.6 | 23.5 | 6.1 | 53.5 | 16.9 | 51.3 | 9.5 | 10.0 | 47.6 | 14.0 |
| InternVL2-40B | 24.5 | 30.8 | 57.4 | 73.2 | 49.2 | 17.1 | 68.3 | 32.7 | 37.8 | 35.3 | 21.2 | 72.1 | 28.6 | 74.4 | 28.6 | 15.0 | 69.0 | 41.9 |
| XgenMM | 0.0 | 0.0 | 0.0 | 1.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Qwen2-VL | 9.4 | 23.1 | 16.7 | 40.2 | 20.6 | 2.9 | 17.1 | 4.1 | 18.4 | 23.5 | 27.3 | 14.0 | 13.0 | 43.6 | 4.8 | 15.0 | 26.2 | 18.6 |
| Qwen2.5-VL | 35.8 | 41.0 | 61.1 | 69.1 | 54.0 | 28.6 | 51.2 | 24.5 | 53.1 | 60.8 | 57.6 | 30.2 | 26.0 | 64.1 | 16.7 | 35.0 | 59.5 | 30.2 |
| GeminiPro | 18.9 | 20.5 | 25.9 | 62.9 | 33.3 | 11.4 | 48.8 | 22.4 | 30.6 | 27.5 | 36.4 | 23.3 | 9.1 | 25.6 | 11.9 | 0.0 | 42.9 | 7.0 |
| Gemini1.5Pro | 45.3 | 46.2 | 64.8 | 76.3 | 57.1 | 28.6 | 80.5 | 44.9 | 75.5 | 62.7 | 66.7 | 79.1 | 31.2 | 87.2 | 38.1 | 20.0 | 57.1 | 32.6 |
| GPT4V | 69.8 | 41.0 | 72.2 | 46.4 | 61.9 | 17.1 | 82.9 | 57.1 | 80.6 | 78.4 | 81.8 | 86.0 | 37.7 | 84.6 | 26.2 | 20.0 | 54.8 | 60.5 |
| GPT4o-mini | 47.2 | 33.3 | 50.0 | 41.2 | 42.9 | 14.3 | 75.6 | 46.9 | 63.3 | 49.0 | 66.7 | 69.8 | 32.5 | 74.4 | 19.0 | 20.0 | 33.3 | 39.5 |
| GPT4o | 60.4 | 56.4 | 68.5 | 43.3 | 57.1 | 40.0 | 80.5 | 46.9 | 67.3 | 58.8 | 63.6 | 65.1 | 37.7 | 79.5 | 40.5 | 25.0 | 71.4 | 46.5 |
| UPD Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 1.9 | 17.9 | 3.7 | 1.0 | 34.9 | 0.0 | 4.9 | 2.0 | 1.0 | 0.0 | 3.0 | 32.6 | 3.9 | 12.8 | 9.5 | 15.0 | 23.8 | 9.3 |
| LLaVA-NeXT-13B | 28.3 | 30.8 | 38.9 | 48.5 | 57.1 | 62.9 | 43.9 | 49.0 | 9.2 | 29.4 | 6.1 | 67.4 | 23.4 | 53.8 | 59.5 | 30.0 | 78.6 | 51.2 |
| LLaVA-NeXT-34B | 62.3 | 76.9 | 63.0 | 71.1 | 60.3 | 80.0 | 90.2 | 67.3 | 51.0 | 78.4 | 45.5 | 88.4 | 40.3 | 66.7 | 66.7 | 30.0 | 90.5 | 62.8 |
| LLaVA-OV-0.5B | 9.4 | 23.1 | 27.8 | 29.9 | 50.8 | 14.3 | 31.7 | 28.6 | 28.6 | 27.5 | 12.1 | 11.6 | 20.8 | 38.5 | 9.5 | 25.0 | 50.0 | 2.3 |
| LLaVA-OV-7B | 0.0 | 2.6 | 1.9 | 27.8 | 1.6 | 2.9 | 0.0 | 4.1 | 2.0 | 5.9 | 3.0 | 2.3 | 0.0 | 10.3 | 4.8 | 0.0 | 9.5 | 4.7 |
| CogVLM-17B | 0.0 | 5.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.3 | 2.6 | 0.0 | 0.0 | 2.4 | 2.3 |
| CogVLM2-19B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics2-8B | 0.0 | 0.0 | 1.9 | 4.1 | 0.0 | 2.9 | 0.0 | 4.1 | 1.0 | 0.0 | 3.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics3-8B | 0.0 | 2.6 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.3 | 0.0 | 7.1 | 0.0 | 2.4 | 0.0 |
| Phi3V | 0.0 | 0.0 | 1.9 | 0.0 | 1.6 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Phi3.5V | 0.0 | 0.0 | 1.9 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.0 | 0.0 | 3.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| InternVL2-2B | 1.9 | 25.6 | 37.0 | 16.5 | 34.9 | 0.0 | 14.6 | 32.7 | 3.1 | 7.8 | 12.1 | 16.3 | 9.1 | 48.7 | 11.9 | 30.0 | 38.1 | 18.6 |
| InternVL2-8B | 20.8 | 38.5 | 44.4 | 44.3 | 44.4 | 42.9 | 39.0 | 32.7 | 32.7 | 27.5 | 6.1 | 58.1 | 23.4 | 53.8 | 16.7 | 15.0 | 47.6 | 27.9 |
| InternVL2-40B | 26.4 | 43.6 | 59.3 | 76.3 | 54.0 | 42.9 | 70.7 | 38.8 | 39.8 | 39.2 | 24.2 | 76.7 | 35.1 | 82.1 | 38.1 | 25.0 | 78.6 | 60.5 |
| XgenMM | 0.0 | 0.0 | 0.0 | 1.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Qwen2-VL | 9.4 | 35.9 | 18.5 | 40.2 | 23.8 | 22.9 | 17.1 | 8.2 | 18.4 | 25.5 | 27.3 | 14.0 | 15.6 | 51.3 | 21.4 | 25.0 | 26.2 | 27.9 |
| Qwen2.5-VL | 35.8 | 56.4 | 63.0 | 71.1 | 58.7 | 65.7 | 53.7 | 34.7 | 54.1 | 66.7 | 60.6 | 41.9 | 37.7 | 64.1 | 47.6 | 55.0 | 69.0 | 39.5 |
| GeminiPro | 20.8 | 61.5 | 29.6 | 66.0 | 39.7 | 20.0 | 51.2 | 26.5 | 34.7 | 33.3 | 36.4 | 27.9 | 24.7 | 38.5 | 26.2 | 15.0 | 54.8 | 18.6 |
| Gemini1.5Pro | 64.2 | 76.9 | 74.1 | 81.4 | 73.0 | 77.1 | 90.2 | 57.1 | 81.6 | 66.7 | 75.8 | 86.0 | 77.9 | 89.7 | 81.0 | 75.0 | 83.3 | 53.5 |
| GPT4V | 79.2 | 94.9 | 81.5 | 80.4 | 77.8 | 88.6 | 87.8 | 77.6 | 82.7 | 82.4 | 87.9 | 88.4 | 77.9 | 92.3 | 92.9 | 85.0 | 95.2 | 86.0 |
| GPT4o-mini | 50.9 | 89.7 | 57.4 | 82.5 | 55.6 | 65.7 | 75.6 | 57.1 | 66.3 | 54.9 | 72.7 | 76.7 | 70.1 | 82.1 | 59.5 | 70.0 | 78.6 | 58.1 |
| GPT4o | 60.4 | 79.5 | 68.5 | 82.5 | 63.5 | 68.6 | 80.5 | 61.2 | 67.3 | 60.8 | 66.7 | 65.1 | 57.1 | 84.6 | 76.2 | 80.0 | 85.7 | 65.1 |
| Standard Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 90.6 | 59.0 | 88.9 | 80.4 | 74.6 | 42.9 | 95.1 | 83.7 | 94.9 | 68.6 | 84.8 | 41.9 | 45.5 | 61.5 | 26.2 | 35.0 | 88.1 | 16.3 |
| LLaVA-NeXT-13B | 88.7 | 59.0 | 81.5 | 83.5 | 66.7 | 31.4 | 100.0 | 75.5 | 95.9 | 84.3 | 87.9 | 55.8 | 49.4 | 69.2 | 19.0 | 50.0 | 59.5 | 20.9 |
| LLaVA-NeXT-34B | 69.8 | 56.4 | 72.2 | 77.3 | 65.1 | 17.1 | 85.4 | 61.2 | 95.9 | 72.5 | 97.0 | 83.7 | 58.4 | 76.9 | 33.3 | 40.0 | 61.9 | 53.5 |
| LLaVA-OV-0.5B | 88.7 | 5.1 | 63.0 | 58.8 | 61.9 | 11.4 | 92.7 | 73.5 | 94.9 | 47.1 | 84.8 | 37.2 | 33.8 | 61.5 | 2.4 | 35.0 | 73.8 | 7.0 |
| LLaVA-OV-7B | 100.0 | 46.2 | 96.3 | 86.6 | 87.3 | 60.0 | 97.6 | 81.6 | 98.0 | 90.2 | 97.0 | 83.7 | 63.6 | 89.7 | 52.4 | 75.0 | 92.9 | 44.2 |
| CogVLM-17B | 86.8 | 41.0 | 74.1 | 85.6 | 77.8 | 14.3 | 97.6 | 75.5 | 95.9 | 84.3 | 75.8 | 58.1 | 29.9 | 71.8 | 28.6 | 20.0 | 88.1 | 7.0 |
| CogVLM2-19B | 94.3 | 66.7 | 90.7 | 92.8 | 85.7 | 42.9 | 100.0 | 83.7 | 96.9 | 90.2 | 100.0 | 79.1 | 57.1 | 82.1 | 54.8 | 70.0 | 85.7 | 30.2 |
| idefics2-8B | 96.2 | 64.1 | 94.4 | 88.7 | 82.5 | 51.4 | 100.0 | 75.5 | 95.9 | 84.3 | 87.9 | 79.1 | 48.1 | 61.5 | 42.9 | 65.0 | 88.1 | 27.9 |
| idefics3-8B | 90.6 | 64.1 | 90.7 | 83.5 | 82.5 | 45.7 | 92.7 | 79.6 | 95.9 | 84.3 | 100.0 | 74.4 | 49.4 | 69.2 | 45.2 | 65.0 | 88.1 | 32.6 |
| Phi3V | 83.0 | 74.4 | 88.9 | 81.4 | 81.0 | 31.4 | 100.0 | 83.7 | 96.9 | 86.3 | 93.9 | 60.5 | 59.7 | 79.5 | 54.8 | 80.0 | 85.7 | 46.5 |
| Phi3.5V | 83.0 | 74.4 | 87.0 | 79.4 | 79.4 | 40.0 | 100.0 | 79.6 | 99.0 | 92.2 | 87.9 | 69.8 | 51.9 | 74.4 | 52.4 | 60.0 | 88.1 | 51.2 |
| InternVL2-2B | 94.3 | 35.9 | 87.0 | 75.3 | 66.7 | 34.3 | 92.7 | 81.6 | 98.0 | 78.4 | 87.9 | 32.6 | 57.1 | 87.2 | 31.0 | 80.0 | 90.5 | 32.6 |
| InternVL2-8B | 90.6 | 74.4 | 92.6 | 85.6 | 74.6 | 40.0 | 95.1 | 83.7 | 96.9 | 82.4 | 87.9 | 69.8 | 54.5 | 87.2 | 50.0 | 75.0 | 90.5 | 53.5 |
| InternVL2-40B | 92.5 | 71.8 | 94.4 | 93.8 | 87.3 | 40.0 | 97.6 | 83.7 | 95.9 | 88.2 | 87.9 | 95.3 | 64.9 | 92.3 | 59.5 | 70.0 | 85.7 | 60.5 |
| XgenMM | 90.6 | 76.9 | 94.4 | 90.7 | 87.3 | 48.6 | 97.6 | 83.7 | 95.9 | 90.2 | 90.9 | 60.5 | 64.9 | 76.9 | 35.7 | 70.0 | 90.5 | 27.9 |
| Qwen2-VL | 90.6 | 66.7 | 92.6 | 94.8 | 79.4 | 37.1 | 100.0 | 77.6 | 95.9 | 86.3 | 90.9 | 76.7 | 57.1 | 84.6 | 45.2 | 45.0 | 90.5 | 53.5 |
| Qwen2.5-VL | 90.6 | 71.8 | 92.6 | 95.9 | 85.7 | 48.6 | 97.6 | 79.6 | 96.9 | 90.2 | 90.9 | 74.4 | 62.3 | 97.4 | 50.0 | 45.0 | 88.1 | 74.4 |
| GeminiPro | 86.8 | 25.6 | 72.2 | 86.6 | 77.8 | 28.6 | 95.1 | 75.5 | 92.9 | 68.6 | 87.9 | 62.8 | 24.7 | 71.8 | 38.1 | 15.0 | 71.4 | 46.5 |
| Gemini1.5Pro | 73.6 | 56.4 | 77.8 | 90.7 | 79.4 | 37.1 | 90.2 | 73.5 | 90.8 | 82.4 | 81.8 | 90.7 | 33.8 | 97.4 | 42.9 | 35.0 | 64.3 | 69.8 |
| GPT4V | 88.7 | 46.2 | 81.5 | 55.7 | 81.0 | 25.7 | 95.1 | 79.6 | 96.9 | 92.2 | 93.9 | 95.3 | 39.0 | 89.7 | 26.2 | 20.0 | 59.5 | 69.8 |
| GPT4o-mini | 92.5 | 38.5 | 79.6 | 55.7 | 85.7 | 25.7 | 97.6 | 77.6 | 95.9 | 84.3 | 87.9 | 86.0 | 36.4 | 92.3 | 35.7 | 25.0 | 45.2 | 72.1 |
| GPT4o | 94.3 | 64.1 | 88.9 | 56.7 | 87.3 | 54.3 | 100.0 | 75.5 | 99.0 | 84.3 | 90.9 | 100.0 | 48.1 | 94.9 | 59.5 | 30.0 | 83.3 | 76.7 |
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #10 | #11 | #12 | #13 | #14 | #15 | #16 | #17 | #18 | |
| Dual Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 35.8 | 41.0 | 64.8 | 58.8 | 49.2 | 25.7 | 78.0 | 59.2 | 57.1 | 52.9 | 42.4 | 39.5 | 32.5 | 41.0 | 19.0 | 20.0 | 61.9 | 4.7 |
| LLaVA-NeXT-13B | 28.3 | 12.8 | 20.4 | 34.0 | 25.4 | 17.1 | 43.9 | 46.9 | 37.8 | 43.1 | 33.3 | 27.9 | 19.5 | 33.3 | 14.3 | 20.0 | 54.8 | 4.7 |
| LLaVA-NeXT-34B | 32.1 | 10.3 | 24.1 | 16.5 | 34.9 | 14.3 | 12.2 | 12.2 | 38.8 | 33.3 | 36.4 | 37.2 | 13.0 | 20.5 | 21.4 | 15.0 | 7.1 | 9.3 |
| LLaVA-OV-0.5B | 22.6 | 0.0 | 7.4 | 11.3 | 15.9 | 8.6 | 14.6 | 2.0 | 26.5 | 9.8 | 33.3 | 0.0 | 10.4 | 2.6 | 0.0 | 10.0 | 14.3 | 0.0 |
| LLaVA-OV-7B | 58.5 | 12.8 | 40.7 | 41.2 | 31.7 | 28.6 | 51.2 | 28.6 | 44.9 | 49.0 | 33.3 | 41.9 | 23.4 | 30.8 | 28.6 | 30.0 | 47.6 | 25.6 |
| CogVLM | 20.8 | 2.6 | 18.5 | 22.7 | 25.4 | 20.0 | 22.0 | 28.6 | 18.4 | 25.5 | 21.2 | 16.3 | 5.2 | 17.9 | 2.4 | 5.0 | 47.6 | 0.0 |
| CogVLM2-19B | 52.8 | 48.7 | 59.3 | 59.8 | 60.3 | 22.9 | 61.0 | 49.0 | 57.1 | 58.8 | 78.8 | 65.1 | 37.7 | 51.3 | 23.8 | 35.0 | 73.8 | 11.6 |
| idefics2-8B | 60.4 | 41.0 | 59.3 | 28.9 | 52.4 | 20.0 | 46.3 | 34.7 | 52.0 | 51.0 | 45.5 | 46.5 | 18.2 | 30.8 | 7.1 | 35.0 | 69.0 | 7.0 |
| idefics3-8B | 58.5 | 61.5 | 61.1 | 50.5 | 61.9 | 37.1 | 46.3 | 49.0 | 52.0 | 56.9 | 51.5 | 58.1 | 39.0 | 46.2 | 26.2 | 60.0 | 73.8 | 18.6 |
| Phi3V | 45.3 | 46.2 | 55.6 | 49.5 | 55.6 | 17.1 | 46.3 | 44.9 | 61.2 | 52.9 | 57.6 | 51.2 | 32.5 | 53.8 | 21.4 | 45.0 | 59.5 | 18.6 |
| Phi3.5V | 22.6 | 41.0 | 31.5 | 13.4 | 22.2 | 11.4 | 29.3 | 14.3 | 34.7 | 27.5 | 24.2 | 32.6 | 11.7 | 25.6 | 23.8 | 10.0 | 28.6 | 27.9 |
| InternVLM2-2B | 50.9 | 33.3 | 72.2 | 53.6 | 50.8 | 22.9 | 70.7 | 46.9 | 73.5 | 45.1 | 57.6 | 30.2 | 40.3 | 43.6 | 33.3 | 75.0 | 71.4 | 18.6 |
| InternVLM2-8B | 67.9 | 43.6 | 81.5 | 82.5 | 73.0 | 45.7 | 82.9 | 49.0 | 82.7 | 74.5 | 75.8 | 67.4 | 59.7 | 53.8 | 26.2 | 60.0 | 73.8 | 41.9 |
| InternVLM2-40B | 69.8 | 61.5 | 70.4 | 73.2 | 60.3 | 40.0 | 61.0 | 53.1 | 69.4 | 64.7 | 57.6 | 65.1 | 42.9 | 61.5 | 28.6 | 55.0 | 76.2 | 39.5 |
| XgenMM | 49.1 | 79.5 | 57.4 | 58.8 | 58.7 | 28.6 | 61.0 | 46.9 | 63.3 | 62.7 | 66.7 | 53.5 | 39.0 | 56.4 | 19.0 | 45.0 | 57.1 | 16.3 |
| Qwen2-VL | 60.4 | 53.8 | 51.9 | 52.6 | 44.4 | 28.6 | 56.1 | 42.9 | 58.2 | 60.8 | 54.5 | 58.1 | 32.5 | 53.8 | 28.6 | 45.0 | 73.8 | 37.2 |
| Qwen2.5-VL | 67.9 | 76.9 | 75.9 | 81.4 | 76.2 | 48.6 | 70.7 | 71.4 | 80.6 | 82.4 | 75.8 | 58.1 | 51.9 | 74.4 | 42.9 | 45.0 | 88.1 | 55.8 |
| GeminiPro | 54.7 | 25.6 | 51.9 | 64.9 | 52.4 | 28.6 | 65.9 | 51.0 | 70.4 | 64.7 | 66.7 | 46.5 | 14.3 | 53.8 | 26.2 | 10.0 | 57.1 | 18.6 |
| Gemini1.5Pro | 60.4 | 41.0 | 63.0 | 82.5 | 74.6 | 40.0 | 85.4 | 73.5 | 82.7 | 70.6 | 57.6 | 81.4 | 29.9 | 94.9 | 35.7 | 25.0 | 69.0 | 72.1 |
| GPT4V | 79.2 | 30.8 | 64.8 | 74.2 | 73.0 | 20.0 | 87.8 | 63.3 | 92.9 | 80.4 | 93.9 | 76.7 | 26.0 | 84.6 | 26.2 | 25.0 | 69.0 | 65.1 |
| GPT4o-mini | 75.5 | 48.7 | 61.1 | 66.0 | 77.8 | 34.3 | 78.0 | 67.3 | 82.7 | 78.4 | 81.8 | 69.8 | 24.7 | 79.5 | 28.6 | 15.0 | 61.9 | 65.1 |
| GPT4o | 79.2 | 61.5 | 75.9 | 52.6 | 76.2 | 45.7 | 82.9 | 55.1 | 84.7 | 80.4 | 78.8 | 86.0 | 41.6 | 84.6 | 59.5 | 30.0 | 76.2 | 81.4 |
| UPD Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 58.5 | 66.7 | 72.2 | 74.2 | 65.1 | 42.9 | 78.0 | 73.5 | 60.2 | 66.7 | 48.5 | 86.0 | 63.6 | 66.7 | 71.4 | 65.0 | 73.8 | 62.8 |
| LLaVA-NeXT-13B | 35.8 | 15.4 | 24.1 | 43.3 | 34.9 | 28.6 | 43.9 | 59.2 | 38.8 | 54.9 | 33.3 | 48.8 | 31.2 | 35.9 | 42.9 | 40.0 | 57.1 | 30.2 |
| LLaVA-NeXT-34B | 35.8 | 12.8 | 29.6 | 20.6 | 44.4 | 22.9 | 14.6 | 16.3 | 39.8 | 33.3 | 36.4 | 46.5 | 18.2 | 25.6 | 47.6 | 35.0 | 7.1 | 18.6 |
| LLaVA-OV-0.5B | 22.6 | 5.1 | 9.3 | 17.5 | 22.2 | 17.1 | 14.6 | 6.1 | 26.5 | 13.7 | 33.3 | 2.3 | 19.5 | 2.6 | 7.1 | 15.0 | 14.3 | 0.0 |
| LLaVA-OV-7B | 60.4 | 53.8 | 42.6 | 50.5 | 39.7 | 48.6 | 51.2 | 40.8 | 45.9 | 52.9 | 36.4 | 51.2 | 42.9 | 30.8 | 61.9 | 55.0 | 50.0 | 39.5 |
| CogVLM | 20.8 | 5.1 | 25.9 | 23.7 | 38.1 | 31.4 | 22.0 | 34.7 | 19.4 | 27.5 | 24.2 | 25.6 | 28.6 | 17.9 | 19.0 | 25.0 | 52.4 | 0.0 |
| CogVLM2-19B | 56.6 | 64.1 | 63.0 | 66.0 | 74.6 | 40.0 | 61.0 | 65.3 | 59.2 | 66.7 | 78.8 | 79.1 | 67.5 | 51.3 | 52.4 | 45.0 | 83.3 | 60.5 |
| idefics2-8B | 67.9 | 69.2 | 66.7 | 33.0 | 68.3 | 34.3 | 51.2 | 51.0 | 57.1 | 66.7 | 63.6 | 67.4 | 58.4 | 48.7 | 54.8 | 60.0 | 81.0 | 44.2 |
| idefics3-8B | 64.2 | 87.2 | 66.7 | 61.9 | 74.6 | 85.7 | 53.7 | 67.3 | 58.2 | 70.6 | 57.6 | 76.7 | 75.3 | 51.3 | 66.7 | 80.0 | 81.0 | 74.4 |
| Phi3V | 56.6 | 64.1 | 61.1 | 63.9 | 68.3 | 34.3 | 46.3 | 59.2 | 64.3 | 68.6 | 63.6 | 81.4 | 61.0 | 64.1 | 52.4 | 60.0 | 66.7 | 86.0 |
| Phi3.5V | 24.5 | 53.8 | 33.3 | 19.6 | 31.7 | 14.3 | 29.3 | 22.4 | 34.7 | 29.4 | 27.3 | 44.2 | 24.7 | 30.8 | 54.8 | 15.0 | 33.3 | 44.2 |
| InternVLM2-2B | 52.8 | 82.1 | 77.8 | 66.0 | 69.8 | 71.4 | 70.7 | 63.3 | 75.5 | 62.7 | 66.7 | 79.1 | 70.1 | 56.4 | 71.4 | 95.0 | 81.0 | 79.1 |
| InternVLM2-8B | 83.0 | 76.9 | 85.2 | 88.7 | 87.3 | 77.1 | 82.9 | 65.3 | 85.7 | 86.3 | 84.8 | 83.7 | 87.0 | 61.5 | 52.4 | 90.0 | 78.6 | 76.7 |
| InternVLM2-40B | 73.6 | 76.9 | 72.2 | 76.3 | 71.4 | 65.7 | 61.0 | 69.4 | 72.4 | 70.6 | 63.6 | 79.1 | 53.2 | 64.1 | 54.8 | 70.0 | 76.2 | 67.4 |
| XgenMM | 56.6 | 92.3 | 61.1 | 67.0 | 68.3 | 45.7 | 65.9 | 63.3 | 65.3 | 70.6 | 72.7 | 83.7 | 67.5 | 64.1 | 69.0 | 75.0 | 66.7 | 67.4 |
| Qwen2-VL | 66.0 | 69.2 | 57.4 | 56.7 | 55.6 | 57.1 | 56.1 | 53.1 | 61.2 | 64.7 | 63.6 | 79.1 | 50.6 | 59.0 | 59.5 | 75.0 | 76.2 | 74.4 |
| Qwen2.5-VL | 81.1 | 92.3 | 77.8 | 85.6 | 88.9 | 88.6 | 73.2 | 87.8 | 82.7 | 90.2 | 81.8 | 90.7 | 84.4 | 84.6 | 92.9 | 95.0 | 92.9 | 90.7 |
| GeminiPro | 60.4 | 59.0 | 70.4 | 73.2 | 74.6 | 51.4 | 68.3 | 73.5 | 73.5 | 86.3 | 81.8 | 69.8 | 53.2 | 59.0 | 59.5 | 50.0 | 71.4 | 32.6 |
| Gemini1.5Pro | 88.7 | 84.6 | 92.6 | 92.8 | 98.4 | 94.3 | 92.7 | 91.8 | 96.9 | 94.1 | 87.9 | 93.0 | 97.4 | 97.4 | 90.5 | 95.0 | 97.6 | 93.0 |
| GPT4V | 96.2 | 97.4 | 90.7 | 92.8 | 90.5 | 91.4 | 92.7 | 91.8 | 95.9 | 96.1 | 97.0 | 90.7 | 93.5 | 94.9 | 95.2 | 100.0 | 95.2 | 95.3 |
| GPT4o-mini | 90.6 | 92.3 | 85.2 | 88.7 | 96.8 | 85.7 | 78.0 | 89.8 | 88.8 | 94.1 | 97.0 | 88.4 | 89.6 | 87.2 | 83.3 | 95.0 | 97.6 | 93.0 |
| GPT4o | 86.8 | 94.9 | 90.7 | 93.8 | 92.1 | 82.9 | 85.4 | 83.7 | 87.8 | 90.2 | 90.9 | 88.4 | 87.0 | 84.6 | 95.2 | 95.0 | 92.9 | 90.7 |
| Standard Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 64.2 | 61.5 | 88.9 | 77.3 | 73.0 | 45.7 | 92.7 | 85.7 | 94.9 | 72.5 | 84.8 | 48.8 | 44.2 | 59.0 | 21.4 | 35.0 | 88.1 | 11.6 |
| LLaVA-NeXT-13B | 83.0 | 61.5 | 83.3 | 81.4 | 74.6 | 48.6 | 97.6 | 81.6 | 95.9 | 86.3 | 90.9 | 60.5 | 48.1 | 66.7 | 31.0 | 50.0 | 92.9 | 25.6 |
| LLaVA-NeXT-34B | 86.8 | 74.4 | 87.0 | 84.5 | 82.5 | 54.3 | 97.6 | 83.7 | 96.9 | 94.1 | 93.9 | 86.0 | 57.1 | 79.5 | 45.2 | 45.0 | 95.2 | 58.1 |
| LLaVA-OV-0.5B | 88.7 | 0.0 | 77.8 | 70.1 | 74.6 | 14.3 | 97.6 | 83.7 | 98.0 | 70.6 | 90.9 | 39.5 | 44.2 | 69.2 | 9.5 | 35.0 | 88.1 | 7.0 |
| LLaVA-OV-7B | 98.1 | 43.6 | 96.3 | 88.7 | 87.3 | 54.3 | 97.6 | 81.6 | 98.0 | 90.2 | 97.0 | 81.4 | 62.3 | 92.3 | 52.4 | 75.0 | 92.9 | 55.8 |
| CogVLM | 88.7 | 56.4 | 77.8 | 85.6 | 76.2 | 37.1 | 97.6 | 77.6 | 93.9 | 84.3 | 75.8 | 53.5 | 16.9 | 71.8 | 19.0 | 15.0 | 88.1 | 4.7 |
| CogVLM2-19B | 94.3 | 69.2 | 92.6 | 90.7 | 85.7 | 42.9 | 100.0 | 81.6 | 95.9 | 90.2 | 100.0 | 79.1 | 53.2 | 82.1 | 54.8 | 70.0 | 85.7 | 30.2 |
| idefics2-8B | 86.8 | 53.8 | 87.0 | 85.6 | 81.0 | 51.4 | 90.2 | 77.6 | 90.8 | 80.4 | 75.8 | 74.4 | 44.2 | 56.4 | 19.0 | 55.0 | 83.3 | 27.9 |
| idefics3-8B | 90.6 | 66.7 | 92.6 | 83.5 | 87.3 | 42.9 | 90.2 | 77.6 | 89.8 | 80.4 | 87.9 | 74.4 | 50.6 | 71.8 | 42.9 | 75.0 | 92.9 | 27.9 |
| Phi3V | 77.4 | 71.8 | 88.9 | 81.4 | 79.4 | 28.6 | 100.0 | 81.6 | 95.9 | 82.4 | 90.9 | 58.1 | 57.1 | 76.9 | 50.0 | 80.0 | 88.1 | 27.9 |
| Phi3.5V | 83.0 | 74.4 | 90.7 | 75.3 | 79.4 | 48.6 | 100.0 | 81.6 | 99.0 | 92.2 | 87.9 | 62.8 | 51.9 | 74.4 | 47.6 | 65.0 | 88.1 | 51.2 |
| InternVLM2-2B | 94.3 | 38.5 | 92.6 | 81.4 | 71.4 | 37.1 | 97.6 | 81.6 | 98.0 | 74.5 | 87.9 | 41.9 | 61.0 | 76.9 | 52.4 | 75.0 | 90.5 | 30.2 |
| InternVLM2-8B | 81.1 | 61.5 | 94.4 | 91.8 | 85.7 | 54.3 | 97.6 | 83.7 | 95.9 | 86.3 | 90.9 | 74.4 | 70.1 | 87.2 | 54.8 | 70.0 | 90.5 | 58.1 |
| InternVLM2-40B | 94.3 | 82.1 | 96.3 | 96.9 | 88.9 | 60.0 | 100.0 | 81.6 | 96.9 | 90.2 | 93.9 | 86.0 | 76.6 | 94.9 | 61.9 | 80.0 | 95.2 | 65.1 |
| XgenMM | 84.9 | 84.6 | 90.7 | 88.7 | 87.3 | 45.7 | 92.7 | 83.7 | 95.9 | 88.2 | 84.8 | 58.1 | 64.9 | 79.5 | 45.2 | 65.0 | 88.1 | 27.9 |
| Qwen2-VL | 90.6 | 69.2 | 92.6 | 93.8 | 81.0 | 51.4 | 100.0 | 81.6 | 96.9 | 94.1 | 87.9 | 76.7 | 63.6 | 82.1 | 38.1 | 55.0 | 95.2 | 48.8 |
| Qwen2.5-VL | 86.8 | 82.1 | 96.3 | 95.9 | 87.3 | 54.3 | 97.6 | 83.7 | 98.0 | 92.2 | 93.9 | 67.4 | 61.0 | 89.7 | 47.6 | 50.0 | 95.2 | 62.8 |
| GeminiPro | 90.6 | 35.9 | 64.8 | 88.7 | 74.6 | 37.1 | 97.6 | 71.4 | 92.9 | 72.5 | 84.8 | 62.8 | 28.6 | 76.9 | 40.5 | 20.0 | 83.3 | 53.5 |
| Gemini1.5Pro | 69.8 | 48.7 | 68.5 | 89.7 | 76.2 | 42.9 | 92.7 | 77.6 | 84.7 | 74.5 | 66.7 | 88.4 | 31.2 | 94.9 | 38.1 | 25.0 | 71.4 | 74.4 |
| GPT4V | 81.1 | 33.3 | 70.4 | 80.4 | 81.0 | 25.7 | 95.1 | 71.4 | 95.9 | 84.3 | 97.0 | 83.7 | 27.3 | 89.7 | 28.6 | 25.0 | 73.8 | 65.1 |
| GPT4o-mini | 83.0 | 53.8 | 72.2 | 77.3 | 81.0 | 40.0 | 100.0 | 77.6 | 92.9 | 84.3 | 84.8 | 76.7 | 28.6 | 92.3 | 31.0 | 20.0 | 64.3 | 65.1 |
| GPT4o | 92.5 | 64.1 | 83.3 | 57.7 | 82.5 | 51.4 | 97.6 | 69.4 | 96.9 | 88.2 | 87.9 | 95.3 | 44.2 | 97.4 | 61.9 | 35.0 | 83.3 | 86.0 |
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #10 | #11 | #12 | #13 | #14 | #15 | #16 | #17 | #18 | |
| Dual Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 39.6 | 51.3 | 68.5 | 57.7 | 63.5 | 31.4 | 78.0 | 65.3 | 69.4 | 56.9 | 60.6 | 41.9 | 33.8 | 51.3 | 19.0 | 20.0 | 78.6 | 7.0 |
| LLaVA-NeXT-13B | 62.3 | 56.4 | 59.3 | 64.9 | 58.7 | 22.9 | 78.0 | 63.3 | 80.6 | 78.4 | 69.7 | 53.5 | 39.0 | 59.0 | 16.7 | 35.0 | 45.2 | 11.6 |
| LLaVA-NeXT-34B | 45.3 | 46.2 | 81.5 | 59.8 | 71.4 | 17.1 | 75.6 | 73.5 | 82.7 | 80.4 | 66.7 | 74.4 | 51.9 | 71.8 | 16.7 | 45.0 | 88.1 | 20.9 |
| LLaVA-OV-0.5B | 1.9 | 0.0 | 7.4 | 8.2 | 7.9 | 0.0 | 0.0 | 6.1 | 5.1 | 7.8 | 6.1 | 0.0 | 0.0 | 7.7 | 0.0 | 0.0 | 0.0 | 0.0 |
| LLaVA-OV-7B | 41.5 | 12.8 | 14.8 | 32.0 | 27.0 | 28.6 | 12.2 | 16.3 | 45.9 | 23.5 | 24.2 | 39.5 | 9.1 | 25.6 | 35.7 | 10.0 | 38.1 | 25.6 |
| CogVLM | 0.0 | 0.0 | 11.1 | 7.2 | 7.9 | 0.0 | 0.0 | 0.0 | 9.2 | 5.9 | 6.1 | 7.0 | 0.0 | 12.8 | 0.0 | 0.0 | 0.0 | 0.0 |
| CogVLM2-19B | 50.9 | 43.6 | 61.1 | 69.1 | 71.4 | 28.6 | 70.7 | 59.2 | 72.4 | 74.5 | 72.7 | 72.1 | 39.0 | 64.1 | 23.8 | 50.0 | 81.0 | 11.6 |
| idefics2-8B | 47.2 | 46.2 | 66.7 | 40.2 | 61.9 | 25.7 | 56.1 | 49.0 | 54.1 | 60.8 | 54.5 | 55.8 | 27.3 | 33.3 | 11.9 | 30.0 | 61.9 | 11.6 |
| idefics3-8B | 66.0 | 61.5 | 68.5 | 50.5 | 63.5 | 37.1 | 58.5 | 53.1 | 60.2 | 60.8 | 72.7 | 44.2 | 42.9 | 48.7 | 21.4 | 40.0 | 61.9 | 9.3 |
| Phi3V | 45.3 | 56.4 | 57.4 | 55.7 | 54.0 | 17.1 | 58.5 | 42.9 | 60.2 | 60.8 | 57.6 | 46.5 | 39.0 | 53.8 | 21.4 | 35.0 | 66.7 | 23.3 |
| Phi3.5V | 37.7 | 51.3 | 35.2 | 30.9 | 36.5 | 22.9 | 31.7 | 22.4 | 45.9 | 47.1 | 39.4 | 34.9 | 13.0 | 25.6 | 21.4 | 10.0 | 50.0 | 27.9 |
| InternVLM2-2B | 1.9 | 15.4 | 25.9 | 26.8 | 41.3 | 17.1 | 12.2 | 4.1 | 12.2 | 11.8 | 15.2 | 20.9 | 15.6 | 33.3 | 9.5 | 10.0 | 21.4 | 14.0 |
| InternVLM2-8B | 64.2 | 48.7 | 72.2 | 69.1 | 61.9 | 45.7 | 68.3 | 42.9 | 67.3 | 54.9 | 60.6 | 60.5 | 45.5 | 46.2 | 35.7 | 45.0 | 73.8 | 18.6 |
| InternVLM2-40B | 77.4 | 69.2 | 88.9 | 94.8 | 74.6 | 8.6 | 95.1 | 73.5 | 94.9 | 76.5 | 84.8 | 79.1 | 64.9 | 89.7 | 42.9 | 60.0 | 78.6 | 48.8 |
| XgenMM | 50.9 | 53.8 | 38.9 | 46.4 | 50.8 | 25.7 | 43.9 | 38.8 | 46.9 | 52.9 | 54.5 | 41.9 | 32.5 | 51.3 | 21.4 | 50.0 | 47.6 | 14.0 |
| Qwen2-VL | 71.7 | 51.3 | 72.2 | 71.1 | 66.7 | 45.7 | 73.2 | 51.0 | 75.5 | 68.6 | 75.8 | 76.7 | 54.5 | 69.2 | 33.3 | 55.0 | 83.3 | 30.2 |
| Qwen2.5-VL | 66.0 | 69.2 | 77.8 | 83.5 | 79.4 | 34.3 | 80.5 | 63.3 | 82.7 | 78.4 | 75.8 | 69.8 | 59.7 | 84.6 | 38.1 | 55.0 | 81.0 | 46.5 |
| GeminiPro | 58.5 | 33.3 | 48.1 | 75.3 | 58.7 | 20.0 | 82.9 | 59.2 | 70.4 | 58.8 | 63.6 | 53.5 | 13.0 | 51.3 | 23.8 | 20.0 | 64.3 | 34.9 |
| Gemini1.5Pro | 45.3 | 35.9 | 63.0 | 83.5 | 73.0 | 40.0 | 90.2 | 71.4 | 71.4 | 68.6 | 51.5 | 86.0 | 24.7 | 87.2 | 26.2 | 10.0 | 61.9 | 46.5 |
| GPT4V | 67.9 | 41.0 | 55.6 | 59.8 | 63.5 | 20.0 | 90.2 | 63.3 | 91.8 | 72.5 | 90.9 | 58.1 | 19.5 | 89.7 | 33.3 | 15.0 | 69.0 | 60.5 |
| GPT4o-mini | 71.7 | 46.2 | 53.7 | 48.5 | 77.8 | 11.4 | 82.9 | 67.3 | 80.6 | 70.6 | 75.8 | 79.1 | 15.6 | 89.7 | 23.8 | 15.0 | 26.2 | 60.5 |
| GPT4o | 79.2 | 61.5 | 75.9 | 48.5 | 81.0 | 28.6 | 85.4 | 65.3 | 88.8 | 76.5 | 81.8 | 86.0 | 31.2 | 94.9 | 52.4 | 30.0 | 73.8 | 76.7 |
| UPD Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 73.6 | 74.4 | 74.1 | 72.2 | 84.1 | 77.1 | 85.4 | 79.6 | 74.5 | 76.5 | 75.8 | 95.3 | 87.0 | 82.1 | 81.0 | 85.0 | 95.2 | 86.0 |
| LLaVA-NeXT-13B | 90.6 | 87.2 | 72.2 | 83.5 | 90.5 | 94.3 | 87.8 | 87.8 | 86.7 | 96.1 | 84.8 | 97.7 | 87.0 | 69.2 | 76.2 | 70.0 | 95.2 | 48.8 |
| LLaVA-NeXT-34B | 94.3 | 100.0 | 100.0 | 97.9 | 98.4 | 97.1 | 97.6 | 98.0 | 94.9 | 100.0 | 97.0 | 97.7 | 94.8 | 94.9 | 97.6 | 100.0 | 97.6 | 97.7 |
| LLaVA-OV-0.5B | 1.9 | 0.0 | 9.3 | 8.2 | 9.5 | 0.0 | 0.0 | 8.2 | 6.1 | 7.8 | 6.1 | 0.0 | 2.6 | 7.7 | 2.4 | 5.0 | 0.0 | 2.3 |
| LLaVA-OV-7B | 41.5 | 48.7 | 14.8 | 38.1 | 33.3 | 54.3 | 12.2 | 28.6 | 46.9 | 25.5 | 24.2 | 44.2 | 13.0 | 25.6 | 64.3 | 25.0 | 42.9 | 62.8 |
| CogVLM | 92.5 | 92.3 | 92.6 | 97.9 | 96.8 | 85.7 | 97.6 | 98.0 | 93.9 | 100.0 | 100.0 | 95.3 | 89.6 | 97.4 | 71.4 | 80.0 | 97.6 | 62.8 |
| CogVLM2-19B | 71.7 | 64.1 | 70.4 | 77.3 | 85.7 | 48.6 | 70.7 | 79.6 | 78.6 | 86.3 | 75.8 | 83.7 | 68.8 | 69.2 | 59.5 | 65.0 | 90.5 | 67.4 |
| idefics2-8B | 71.7 | 74.4 | 70.4 | 45.4 | 77.8 | 54.3 | 58.5 | 65.3 | 55.1 | 72.5 | 69.7 | 74.4 | 67.5 | 41.0 | 45.2 | 55.0 | 76.2 | 37.2 |
| idefics3-8B | 73.6 | 82.1 | 68.5 | 59.8 | 73.0 | 77.1 | 58.5 | 71.4 | 64.3 | 74.5 | 72.7 | 62.8 | 83.1 | 51.3 | 64.3 | 70.0 | 69.0 | 46.5 |
| Phi3V | 64.2 | 74.4 | 63.0 | 68.0 | 69.8 | 68.6 | 58.5 | 63.3 | 63.3 | 72.5 | 66.7 | 83.7 | 66.2 | 66.7 | 64.3 | 55.0 | 78.6 | 81.4 |
| Phi3.5V | 49.1 | 66.7 | 37.0 | 49.5 | 49.2 | 45.7 | 31.7 | 38.8 | 45.9 | 52.9 | 48.5 | 51.2 | 31.2 | 38.5 | 52.4 | 30.0 | 57.1 | 62.8 |
| InternVLM2-2B | 1.9 | 25.6 | 25.9 | 32.0 | 49.2 | 40.0 | 12.2 | 4.1 | 12.2 | 23.5 | 18.2 | 37.2 | 20.8 | 46.2 | 54.8 | 20.0 | 23.8 | 41.9 |
| InternVLM2-8B | 67.9 | 59.0 | 74.1 | 72.2 | 81.0 | 85.7 | 70.7 | 59.2 | 68.4 | 64.7 | 63.6 | 76.7 | 66.2 | 46.2 | 54.8 | 65.0 | 76.2 | 27.9 |
| InternVLM2-40B | 100.0 | 94.9 | 98.1 | 100.0 | 96.8 | 94.3 | 100.0 | 98.0 | 99.0 | 100.0 | 97.0 | 100.0 | 96.1 | 97.4 | 83.3 | 85.0 | 100.0 | 79.1 |
| XgenMM | 54.7 | 69.2 | 40.7 | 50.5 | 60.3 | 54.3 | 43.9 | 55.1 | 49.0 | 56.9 | 57.6 | 65.1 | 53.2 | 64.1 | 61.9 | 85.0 | 54.8 | 58.1 |
| Qwen2-VL | 83.0 | 82.1 | 77.8 | 75.3 | 81.0 | 85.7 | 73.2 | 65.3 | 78.6 | 76.5 | 87.9 | 95.3 | 83.1 | 84.6 | 61.9 | 75.0 | 85.7 | 74.4 |
| Qwen2.5-VL | 83.0 | 94.9 | 88.9 | 89.7 | 93.7 | 94.3 | 87.8 | 79.6 | 85.7 | 88.2 | 81.8 | 90.7 | 88.3 | 89.7 | 85.7 | 95.0 | 95.2 | 90.7 |
| GeminiPro | 69.8 | 87.2 | 74.1 | 86.6 | 74.6 | 71.4 | 87.8 | 75.5 | 75.5 | 92.2 | 84.8 | 79.1 | 64.9 | 71.8 | 71.4 | 60.0 | 83.3 | 58.1 |
| Gemini1.5Pro | 88.7 | 94.9 | 98.1 | 96.9 | 100.0 | 94.3 | 97.6 | 93.9 | 95.9 | 88.2 | 93.9 | 97.7 | 94.8 | 100.0 | 97.6 | 90.0 | 97.6 | 88.4 |
| GPT4V | 100.0 | 97.4 | 96.3 | 97.9 | 98.4 | 100.0 | 97.6 | 95.9 | 98.0 | 98.0 | 100.0 | 100.0 | 94.8 | 100.0 | 95.2 | 100.0 | 97.6 | 97.7 |
| GPT4o-mini | 94.3 | 97.4 | 88.9 | 99.0 | 98.4 | 88.6 | 97.6 | 95.9 | 94.9 | 96.1 | 100.0 | 97.7 | 94.8 | 97.4 | 85.7 | 95.0 | 97.6 | 95.3 |
| GPT4o | 86.8 | 94.9 | 90.7 | 92.8 | 96.8 | 85.7 | 92.7 | 93.9 | 90.8 | 94.1 | 97.0 | 95.3 | 88.3 | 97.4 | 97.6 | 95.0 | 97.6 | 90.7 |
| Standard Acc. | ||||||||||||||||||
| LLaVA1.5-13b | 58.5 | 61.5 | 88.9 | 79.4 | 73.0 | 37.1 | 87.8 | 83.7 | 93.9 | 70.6 | 81.8 | 46.5 | 41.6 | 61.5 | 21.4 | 25.0 | 83.3 | 11.6 |
| LLaVA-NeXT-13B | 67.9 | 59.0 | 75.9 | 79.4 | 63.5 | 22.9 | 87.8 | 75.5 | 92.9 | 82.4 | 75.8 | 55.8 | 44.2 | 66.7 | 16.7 | 35.0 | 47.6 | 16.3 |
| LLaVA-NeXT-34B | 50.9 | 46.2 | 81.5 | 61.9 | 73.0 | 17.1 | 78.0 | 75.5 | 87.8 | 80.4 | 69.7 | 76.7 | 54.5 | 74.4 | 19.0 | 45.0 | 90.5 | 20.9 |
| LLaVA-OV-0.5B | 81.1 | 2.6 | 72.2 | 66.0 | 61.9 | 11.4 | 92.7 | 77.6 | 95.9 | 51.0 | 84.8 | 34.9 | 20.8 | 61.5 | 2.4 | 30.0 | 66.7 | 4.7 |
| LLaVA-OV-7B | 100.0 | 41.0 | 96.3 | 87.6 | 87.3 | 48.6 | 97.6 | 81.6 | 98.0 | 90.2 | 93.9 | 81.4 | 61.0 | 87.2 | 54.8 | 75.0 | 92.9 | 41.9 |
| CogVLM | 0.0 | 0.0 | 13.0 | 7.2 | 9.5 | 0.0 | 0.0 | 0.0 | 12.2 | 5.9 | 6.1 | 7.0 | 0.0 | 12.8 | 0.0 | 0.0 | 0.0 | 0.0 |
| CogVLM2-19B | 71.7 | 64.1 | 87.0 | 88.7 | 84.1 | 42.9 | 95.1 | 75.5 | 90.8 | 88.2 | 97.0 | 83.7 | 55.8 | 79.5 | 52.4 | 70.0 | 85.7 | 25.6 |
| idefics2-8B | 69.8 | 61.5 | 90.7 | 88.7 | 82.5 | 45.7 | 97.6 | 79.6 | 94.9 | 86.3 | 78.8 | 79.1 | 44.2 | 61.5 | 42.9 | 55.0 | 81.0 | 23.3 |
| idefics3-8B | 92.5 | 69.2 | 96.3 | 84.5 | 87.3 | 45.7 | 100.0 | 81.6 | 92.9 | 84.3 | 97.0 | 72.1 | 51.9 | 71.8 | 38.1 | 70.0 | 92.9 | 30.2 |
| Phi3V | 71.7 | 71.8 | 87.0 | 80.4 | 77.8 | 22.9 | 100.0 | 79.6 | 95.9 | 84.3 | 90.9 | 53.5 | 57.1 | 76.9 | 42.9 | 75.0 | 85.7 | 37.2 |
| Phi3.5V | 77.4 | 76.9 | 87.0 | 75.3 | 76.2 | 40.0 | 100.0 | 79.6 | 98.0 | 90.2 | 87.9 | 69.8 | 46.8 | 69.2 | 50.0 | 65.0 | 88.1 | 39.5 |
| InternVLM2-2B | 92.5 | 41.0 | 92.6 | 78.4 | 76.2 | 48.6 | 100.0 | 81.6 | 99.0 | 74.5 | 87.9 | 46.5 | 59.7 | 76.9 | 28.6 | 75.0 | 92.9 | 34.9 |
| InternVLM2-8B | 88.7 | 74.4 | 96.3 | 92.8 | 81.0 | 57.1 | 97.6 | 83.7 | 96.9 | 86.3 | 87.9 | 72.1 | 68.8 | 92.3 | 61.9 | 75.0 | 92.9 | 48.8 |
| InternVLM2-40B | 77.4 | 74.4 | 90.7 | 94.8 | 76.2 | 8.6 | 95.1 | 75.5 | 94.9 | 76.5 | 87.9 | 79.1 | 67.5 | 92.3 | 59.5 | 75.0 | 78.6 | 62.8 |
| XgenMM | 88.7 | 74.4 | 92.6 | 90.7 | 87.3 | 42.9 | 95.1 | 83.7 | 95.9 | 86.3 | 93.9 | 58.1 | 64.9 | 71.8 | 38.1 | 65.0 | 88.1 | 27.9 |
| Qwen2-VL | 86.8 | 66.7 | 92.6 | 91.8 | 85.7 | 57.1 | 100.0 | 79.6 | 96.9 | 90.2 | 87.9 | 81.4 | 62.3 | 79.5 | 54.8 | 55.0 | 95.2 | 39.5 |
| Qwen2.5-VL | 83.0 | 74.4 | 85.2 | 92.8 | 85.7 | 40.0 | 92.7 | 79.6 | 95.9 | 88.2 | 93.9 | 74.4 | 62.3 | 94.9 | 50.0 | 55.0 | 85.7 | 53.5 |
| GeminiPro | 88.7 | 33.3 | 63.0 | 83.5 | 74.6 | 22.9 | 92.7 | 71.4 | 89.8 | 62.7 | 72.7 | 55.8 | 16.9 | 64.1 | 35.7 | 25.0 | 73.8 | 48.8 |
| Gemini1.5Pro | 49.1 | 38.5 | 63.0 | 86.6 | 73.0 | 42.9 | 92.7 | 73.5 | 74.5 | 72.5 | 51.5 | 86.0 | 26.0 | 87.2 | 28.6 | 10.0 | 64.3 | 53.5 |
| GPT4V | 67.9 | 41.0 | 59.3 | 61.9 | 65.1 | 20.0 | 92.7 | 65.3 | 93.9 | 74.5 | 90.9 | 58.1 | 20.8 | 89.7 | 35.7 | 15.0 | 71.4 | 60.5 |
| GPT4o-mini | 77.4 | 46.2 | 63.0 | 49.5 | 79.4 | 17.1 | 85.4 | 71.4 | 85.7 | 70.6 | 75.8 | 79.1 | 16.9 | 92.3 | 31.0 | 20.0 | 28.6 | 62.8 |
| GPT4o | 92.5 | 61.5 | 83.3 | 54.6 | 84.1 | 31.4 | 92.7 | 69.4 | 96.9 | 80.4 | 84.8 | 88.4 | 32.5 | 97.4 | 54.8 | 30.0 | 76.2 | 83.7 |
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #11 | #12 | #17 | |
| Dual Acc. | ||||||||||||
| LLaVA1.5-13b | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| LLaVA-NeXT-13B | 42.9 | 0.0 | 11.1 | 73.5 | 12.5 | 25.0 | 31.1 | 19.4 | 19.4 | 4.3 | 20.0 | 48.8 |
| LLaVA-NeXT-34B | 57.1 | 21.7 | 33.3 | 89.7 | 37.5 | 62.5 | 82.2 | 30.6 | 45.2 | 13.0 | 26.7 | 62.8 |
| LLaVA-OV-0.5B | 7.1 | 0.0 | 16.7 | 5.9 | 0.0 | 12.5 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 53.5 |
| LLaVA-OV-7B | 0.0 | 0.0 | 0.0 | 11.8 | 0.0 | 0.0 | 0.0 | 0.0 | 3.2 | 0.0 | 0.0 | 0.0 |
| CogVLM | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| CogVLM2-19B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics2-8B | 0.0 | 0.0 | 0.0 | 1.5 | 0.0 | 0.0 | 0.0 | 2.8 | 0.0 | 0.0 | 0.0 | 7.0 |
| idefics3-8B | 0.0 | 0.0 | 5.6 | 0.0 | 6.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 25.6 |
| Phi3V | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Phi3.5V | 0.0 | 0.0 | 0.0 | 1.5 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| InternVLM2-2B | 14.3 | 0.0 | 0.0 | 19.1 | 6.2 | 4.2 | 13.3 | 16.7 | 22.6 | 0.0 | 0.0 | 44.2 |
| InternVLM2-8B | 71.4 | 17.4 | 11.1 | 39.7 | 0.0 | 37.5 | 31.1 | 25.0 | 12.9 | 13.0 | 6.7 | 41.9 |
| InternVLM2-40B | 71.4 | 0.0 | 5.6 | 89.7 | 0.0 | 25.0 | 48.9 | 41.7 | 38.7 | 13.0 | 6.7 | 48.8 |
| XgenMM | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Qwen2-VL | 42.9 | 0.0 | 16.7 | 61.8 | 12.5 | 45.8 | 62.2 | 44.4 | 22.6 | 0.0 | 20.0 | 32.6 |
| Qwen2.5-VL | 85.7 | 60.9 | 72.2 | 92.6 | 56.2 | 83.3 | 66.7 | 69.4 | 64.5 | 17.4 | 40.0 | 86.0 |
| GeminiPro | 71.4 | 17.4 | 38.9 | 70.6 | 50.0 | 37.5 | 28.9 | 8.3 | 19.4 | 0.0 | 6.7 | 58.1 |
| Gemini1.5Pro | 78.6 | 60.9 | 44.4 | 95.6 | 62.5 | 87.5 | 91.1 | 27.8 | 80.6 | 43.5 | 26.7 | 62.8 |
| GPT4V | 78.6 | 78.3 | 50.0 | 45.6 | 68.8 | 87.5 | 93.3 | 44.4 | 87.1 | 34.8 | 13.3 | 60.5 |
| GPT4o-mini | 64.3 | 60.9 | 55.6 | 39.7 | 25.0 | 75.0 | 80.0 | 27.8 | 64.5 | 26.1 | 13.3 | 37.2 |
| GPT4o | 100.0 | 78.3 | 66.7 | 39.7 | 56.2 | 87.5 | 97.8 | 47.2 | 64.5 | 56.5 | 20.0 | 79.1 |
| UPD Acc. | ||||||||||||
| LLaVA1.5-13b | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| LLaVA-NeXT-13B | 50.0 | 0.0 | 11.1 | 83.8 | 12.5 | 25.0 | 53.3 | 25.0 | 22.6 | 26.1 | 20.0 | 86.0 |
| LLaVA-NeXT-34B | 85.7 | 21.7 | 50.0 | 98.5 | 37.5 | 66.7 | 95.6 | 69.4 | 45.2 | 43.5 | 53.3 | 97.7 |
| LLaVA-OV-0.5B | 14.3 | 0.0 | 16.7 | 11.8 | 0.0 | 12.5 | 0.0 | 2.8 | 0.0 | 0.0 | 6.7 | 65.1 |
| LLaVA-OV-7B | 0.0 | 0.0 | 0.0 | 14.7 | 0.0 | 0.0 | 0.0 | 0.0 | 3.2 | 0.0 | 0.0 | 0.0 |
| CogVLM | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| CogVLM2-19B | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| idefics2-8B | 0.0 | 0.0 | 0.0 | 1.5 | 0.0 | 0.0 | 0.0 | 2.8 | 0.0 | 0.0 | 0.0 | 9.3 |
| idefics3-8B | 0.0 | 0.0 | 5.6 | 0.0 | 6.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 27.9 |
| Phi3V | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Phi3.5V | 0.0 | 0.0 | 0.0 | 1.5 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| InternVLM2-2B | 14.3 | 0.0 | 0.0 | 23.5 | 12.5 | 8.3 | 26.7 | 22.2 | 22.6 | 8.7 | 0.0 | 44.2 |
| InternVLM2-8B | 71.4 | 17.4 | 11.1 | 48.5 | 0.0 | 37.5 | 51.1 | 41.7 | 12.9 | 13.0 | 13.3 | 46.5 |
| InternVLM2-40B | 71.4 | 0.0 | 5.6 | 94.1 | 0.0 | 29.2 | 51.1 | 52.8 | 38.7 | 21.7 | 20.0 | 62.8 |
| XgenMM | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Qwen2-VL | 50.0 | 0.0 | 22.2 | 63.2 | 12.5 | 50.0 | 75.6 | 63.9 | 25.8 | 4.3 | 40.0 | 37.2 |
| Qwen2.5-VL | 85.7 | 73.9 | 83.3 | 95.6 | 56.2 | 91.7 | 91.1 | 97.2 | 64.5 | 34.8 | 60.0 | 97.7 |
| GeminiPro | 78.6 | 26.1 | 55.6 | 73.5 | 50.0 | 45.8 | 60.0 | 72.2 | 22.6 | 4.3 | 46.7 | 83.7 |
| Gemini1.5Pro | 92.9 | 87.0 | 88.9 | 100.0 | 81.2 | 100.0 | 100.0 | 100.0 | 83.9 | 95.7 | 100.0 | 97.7 |
| GPT4V | 100.0 | 91.3 | 88.9 | 100.0 | 93.8 | 95.8 | 97.8 | 100.0 | 96.8 | 95.7 | 100.0 | 100.0 |
| GPT4o-mini | 71.4 | 78.3 | 88.9 | 98.5 | 31.2 | 87.5 | 95.6 | 94.4 | 74.2 | 65.2 | 86.7 | 97.7 |
| GPT4o | 100.0 | 87.0 | 88.9 | 98.5 | 68.8 | 91.7 | 97.8 | 91.7 | 71.0 | 82.6 | 86.7 | 93.0 |
| Standard Acc. | ||||||||||||
| LLaVA1.5-13b | 92.9 | 82.6 | 77.8 | 88.2 | 68.8 | 87.5 | 44.4 | 55.6 | 54.8 | 13.0 | 26.7 | 88.4 |
| LLaVA-NeXT-13B | 92.9 | 82.6 | 72.2 | 86.8 | 75.0 | 87.5 | 60.0 | 55.6 | 64.5 | 17.4 | 46.7 | 58.1 |
| LLaVA-NeXT-34B | 64.3 | 82.6 | 66.7 | 89.7 | 87.5 | 87.5 | 86.7 | 58.3 | 77.4 | 30.4 | 33.3 | 65.1 |
| LLaVA-OV-0.5B | 92.9 | 13.0 | 50.0 | 58.8 | 75.0 | 79.2 | 35.6 | 33.3 | 54.8 | 0.0 | 40.0 | 76.7 |
| LLaVA-OV-7B | 100.0 | 47.8 | 94.4 | 94.1 | 100.0 | 95.8 | 84.4 | 72.2 | 90.3 | 69.6 | 73.3 | 93.0 |
| CogVLM | 85.7 | 43.5 | 55.6 | 89.7 | 93.8 | 91.7 | 57.8 | 36.1 | 67.7 | 8.7 | 13.3 | 88.4 |
| CogVLM2-19B | 100.0 | 95.7 | 77.8 | 98.5 | 93.8 | 95.8 | 77.8 | 66.7 | 77.4 | 60.9 | 60.0 | 86.0 |
| idefics2-8B | 100.0 | 87.0 | 88.9 | 95.6 | 81.2 | 91.7 | 80.0 | 47.2 | 54.8 | 43.5 | 66.7 | 88.4 |
| idefics3-8B | 92.9 | 73.9 | 88.9 | 94.1 | 87.5 | 91.7 | 73.3 | 50.0 | 64.5 | 56.5 | 66.7 | 88.4 |
| Phi3V | 85.7 | 82.6 | 88.9 | 92.6 | 100.0 | 91.7 | 62.2 | 69.4 | 74.2 | 52.2 | 73.3 | 86.0 |
| Phi3.5V | 71.4 | 87.0 | 66.7 | 88.2 | 87.5 | 91.7 | 68.9 | 58.3 | 71.0 | 47.8 | 53.3 | 88.4 |
| InternVLM2-2B | 100.0 | 60.9 | 88.9 | 75.0 | 81.2 | 87.5 | 35.6 | 75.0 | 83.9 | 34.8 | 80.0 | 90.7 |
| InternVLM2-8B | 100.0 | 82.6 | 94.4 | 89.7 | 81.2 | 83.3 | 68.9 | 61.1 | 87.1 | 78.3 | 80.0 | 90.7 |
| InternVLM2-40B | 92.9 | 91.3 | 94.4 | 95.6 | 100.0 | 91.7 | 95.6 | 75.0 | 93.5 | 91.3 | 80.0 | 86.0 |
| XgenMM | 92.9 | 82.6 | 88.9 | 92.6 | 93.8 | 95.8 | 62.2 | 75.0 | 71.0 | 39.1 | 73.3 | 90.7 |
| Qwen2-VL | 92.9 | 69.6 | 88.9 | 97.1 | 87.5 | 87.5 | 77.8 | 66.7 | 83.9 | 69.6 | 46.7 | 90.7 |
| Qwen2.5-VL | 100.0 | 82.6 | 88.9 | 97.1 | 100.0 | 91.7 | 73.3 | 72.2 | 96.8 | 69.6 | 40.0 | 88.4 |
| GeminiPro | 92.9 | 69.6 | 44.4 | 94.1 | 93.8 | 87.5 | 57.8 | 11.1 | 71.0 | 34.8 | 13.3 | 67.4 |
| Gemini1.5Pro | 85.7 | 65.2 | 50.0 | 95.6 | 75.0 | 87.5 | 91.1 | 27.8 | 93.5 | 47.8 | 26.7 | 65.1 |
| GPT4V | 78.6 | 82.6 | 55.6 | 45.6 | 75.0 | 91.7 | 95.6 | 44.4 | 90.3 | 39.1 | 13.3 | 60.5 |
| GPT4o-mini | 92.9 | 82.6 | 61.1 | 41.2 | 87.5 | 87.5 | 84.4 | 30.6 | 90.3 | 43.5 | 13.3 | 39.5 |
| GPT4o | 100.0 | 87.0 | 77.8 | 41.2 | 87.5 | 95.8 | 100.0 | 52.8 | 93.5 | 73.9 | 20.0 | 83.7 |
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #11 | #12 | #17 | |
| Dual Acc. | ||||||||||||
| LLaVA1.5-13b | 71.4 | 0.0 | 11.1 | 85.3 | 0.0 | 66.7 | 13.3 | 30.6 | 22.6 | 0.0 | 6.7 | 67.4 |
| LLaVA-NeXT-13B | 71.4 | 0.0 | 16.7 | 79.4 | 6.2 | 58.3 | 8.9 | 50.0 | 25.8 | 0.0 | 20.0 | 46.5 |
| LLaVA-NeXT-34B | 57.1 | 21.7 | 16.7 | 92.6 | 12.5 | 62.5 | 46.7 | 41.7 | 35.5 | 4.3 | 26.7 | 74.4 |
| LLaVA-OV-0.5B | 14.3 | 0.0 | 0.0 | 17.6 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 32.6 |
| LLaVA-OV-7B | 92.9 | 0.0 | 38.9 | 89.7 | 12.5 | 66.7 | 35.6 | 41.7 | 32.3 | 17.4 | 20.0 | 76.7 |
| CogVLM | 0.0 | 0.0 | 0.0 | 36.8 | 6.2 | 37.5 | 20.0 | 0.0 | 0.0 | 0.0 | 0.0 | 58.1 |
| CogVLM2-19B | 78.6 | 0.0 | 33.3 | 79.4 | 12.5 | 54.2 | 37.8 | 30.6 | 9.7 | 0.0 | 6.7 | 79.1 |
| idefics2-8B | 71.4 | 4.3 | 27.8 | 88.2 | 31.2 | 70.8 | 62.2 | 16.7 | 19.4 | 0.0 | 13.3 | 81.4 |
| idefics3-8B | 78.6 | 8.7 | 50.0 | 76.5 | 43.8 | 62.5 | 60.0 | 36.1 | 35.5 | 0.0 | 26.7 | 90.7 |
| Phi3V | 50.0 | 17.4 | 33.3 | 89.7 | 31.2 | 83.3 | 51.1 | 44.4 | 54.8 | 0.0 | 33.3 | 83.7 |
| Phi3.5V | 57.1 | 13.0 | 33.3 | 83.8 | 18.8 | 75.0 | 53.3 | 44.4 | 41.9 | 0.0 | 20.0 | 83.7 |
| InternVLM2-2B | 42.9 | 0.0 | 5.6 | 57.4 | 0.0 | 16.7 | 4.4 | 22.2 | 25.8 | 0.0 | 0.0 | 7.0 |
| InternVLM2-8B | 92.9 | 30.4 | 27.8 | 89.7 | 12.5 | 62.5 | 48.9 | 61.1 | 64.5 | 8.7 | 13.3 | 88.4 |
| InternVLM2-40B | 85.7 | 21.7 | 22.2 | 98.5 | 12.5 | 66.7 | 35.6 | 61.1 | 51.6 | 4.3 | 13.3 | 86.0 |
| XgenMM | 85.7 | 39.1 | 33.3 | 85.3 | 37.5 | 75.0 | 51.1 | 61.1 | 38.7 | 8.7 | 13.3 | 86.0 |
| Qwen2-VL | 85.7 | 34.8 | 27.8 | 94.1 | 37.5 | 62.5 | 77.8 | 63.9 | 38.7 | 0.0 | 40.0 | 93.0 |
| Qwen2.5-VL | 100.0 | 56.5 | 72.2 | 97.1 | 68.8 | 87.5 | 68.9 | 63.9 | 71.0 | 30.4 | 40.0 | 90.7 |
| GeminiPro | 92.9 | 52.2 | 50.0 | 85.3 | 62.5 | 79.2 | 55.6 | 11.1 | 45.2 | 8.7 | 6.7 | 86.0 |
| Gemini1.5Pro | 100.0 | 69.6 | 44.4 | 95.6 | 75.0 | 91.7 | 86.7 | 19.4 | 93.5 | 47.8 | 26.7 | 67.4 |
| GPT4V | 85.7 | 73.9 | 50.0 | 69.1 | 81.2 | 91.7 | 57.8 | 19.4 | 83.9 | 30.4 | 13.3 | 69.8 |
| GPT4o-mini | 92.9 | 87.0 | 50.0 | 60.3 | 62.5 | 87.5 | 71.1 | 16.7 | 83.9 | 26.1 | 20.0 | 48.8 |
| GPT4o | 100.0 | 82.6 | 72.2 | 41.2 | 81.2 | 91.7 | 91.1 | 50.0 | 83.9 | 47.8 | 20.0 | 90.7 |
| UPD Acc. | ||||||||||||
| LLaVA1.5-13b | 71.4 | 0.0 | 11.1 | 95.6 | 6.2 | 66.7 | 15.6 | 47.2 | 29.0 | 4.3 | 26.7 | 76.7 |
| LLaVA-NeXT-13B | 85.7 | 0.0 | 16.7 | 92.6 | 6.2 | 62.5 | 13.3 | 83.3 | 29.0 | 0.0 | 26.7 | 48.8 |
| LLaVA-NeXT-34B | 64.3 | 21.7 | 22.2 | 97.1 | 12.5 | 62.5 | 55.6 | 75.0 | 35.5 | 4.3 | 40.0 | 79.1 |
| LLaVA-OV-0.5B | 14.3 | 0.0 | 0.0 | 22.1 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 37.2 |
| LLaVA-OV-7B | 92.9 | 13.0 | 38.9 | 94.1 | 12.5 | 66.7 | 46.7 | 63.9 | 32.3 | 30.4 | 33.3 | 81.4 |
| CogVLM | 0.0 | 0.0 | 0.0 | 41.2 | 6.2 | 37.5 | 42.2 | 0.0 | 0.0 | 0.0 | 0.0 | 60.5 |
| CogVLM2-19B | 78.6 | 0.0 | 33.3 | 80.9 | 12.5 | 58.3 | 44.4 | 52.8 | 9.7 | 0.0 | 20.0 | 90.7 |
| idefics2-8B | 92.9 | 4.3 | 33.3 | 91.2 | 31.2 | 75.0 | 68.9 | 36.1 | 25.8 | 4.3 | 33.3 | 95.3 |
| idefics3-8B | 92.9 | 8.7 | 55.6 | 80.9 | 43.8 | 62.5 | 86.7 | 61.1 | 38.7 | 0.0 | 40.0 | 97.7 |
| Phi3V | 85.7 | 26.1 | 33.3 | 97.1 | 37.5 | 91.7 | 97.8 | 72.2 | 54.8 | 4.3 | 46.7 | 97.7 |
| Phi3.5V | 78.6 | 13.0 | 33.3 | 97.1 | 18.8 | 83.3 | 75.6 | 72.2 | 45.2 | 0.0 | 40.0 | 95.3 |
| InternVLM2-2B | 42.9 | 0.0 | 5.6 | 69.1 | 0.0 | 16.7 | 4.4 | 25.0 | 25.8 | 0.0 | 6.7 | 9.3 |
| InternVLM2-8B | 92.9 | 52.2 | 27.8 | 94.1 | 12.5 | 79.2 | 75.6 | 83.3 | 64.5 | 13.0 | 40.0 | 97.7 |
| InternVLM2-40B | 92.9 | 39.1 | 22.2 | 100.0 | 12.5 | 75.0 | 40.0 | 75.0 | 51.6 | 13.0 | 26.7 | 90.7 |
| XgenMM | 92.9 | 39.1 | 33.3 | 94.1 | 37.5 | 79.2 | 80.0 | 75.0 | 45.2 | 17.4 | 26.7 | 93.0 |
| Qwen2-VL | 92.9 | 47.8 | 33.3 | 98.5 | 37.5 | 75.0 | 95.6 | 88.9 | 45.2 | 4.3 | 53.3 | 97.7 |
| Qwen2.5-VL | 100.0 | 60.9 | 77.8 | 100.0 | 68.8 | 95.8 | 100.0 | 97.2 | 77.4 | 69.6 | 66.7 | 100.0 |
| GeminiPro | 100.0 | 60.9 | 83.3 | 91.2 | 68.8 | 87.5 | 93.3 | 94.4 | 48.4 | 47.8 | 60.0 | 97.7 |
| Gemini1.5Pro | 100.0 | 95.7 | 100.0 | 100.0 | 93.8 | 100.0 | 100.0 | 100.0 | 96.8 | 100.0 | 100.0 | 97.7 |
| GPT4V | 100.0 | 100.0 | 94.4 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 90.3 | 100.0 | 100.0 | 100.0 |
| GPT4o-mini | 100.0 | 100.0 | 100.0 | 100.0 | 81.2 | 100.0 | 100.0 | 100.0 | 93.5 | 82.6 | 93.3 | 100.0 |
| GPT4o | 100.0 | 95.7 | 100.0 | 100.0 | 93.8 | 95.8 | 100.0 | 100.0 | 87.1 | 87.0 | 100.0 | 100.0 |
| Standard Acc. | ||||||||||||
| LLaVA1.5-13b | 85.7 | 82.6 | 83.3 | 88.2 | 68.8 | 91.7 | 51.1 | 61.1 | 51.6 | 8.7 | 26.7 | 90.7 |
| LLaVA-NeXT-13B | 85.7 | 87.0 | 72.2 | 85.3 | 75.0 | 87.5 | 57.8 | 58.3 | 61.3 | 17.4 | 40.0 | 93.0 |
| LLaVA-NeXT-34B | 78.6 | 87.0 | 77.8 | 94.1 | 100.0 | 91.7 | 84.4 | 58.3 | 77.4 | 52.2 | 33.3 | 95.3 |
| LLaVA-OV-0.5B | 92.9 | 8.7 | 55.6 | 76.5 | 93.8 | 87.5 | 42.2 | 36.1 | 67.7 | 0.0 | 40.0 | 83.7 |
| LLaVA-OV-7B | 100.0 | 34.8 | 94.4 | 95.6 | 100.0 | 95.8 | 80.0 | 69.4 | 90.3 | 69.6 | 73.3 | 93.0 |
| CogVLM | 85.7 | 87.0 | 50.0 | 89.7 | 81.2 | 91.7 | 60.0 | 16.7 | 67.7 | 8.7 | 6.7 | 90.7 |
| CogVLM2-19B | 100.0 | 95.7 | 88.9 | 97.1 | 93.8 | 91.7 | 82.2 | 66.7 | 77.4 | 56.5 | 60.0 | 86.0 |
| idefics2-8B | 78.6 | 65.2 | 83.3 | 95.6 | 87.5 | 91.7 | 75.6 | 44.4 | 51.6 | 34.8 | 53.3 | 86.0 |
| idefics3-8B | 85.7 | 78.3 | 94.4 | 94.1 | 93.8 | 91.7 | 73.3 | 55.6 | 67.7 | 60.9 | 73.3 | 93.0 |
| Phi3V | 64.3 | 82.6 | 88.9 | 91.2 | 93.8 | 91.7 | 53.3 | 63.9 | 74.2 | 56.5 | 73.3 | 86.0 |
| Phi3.5V | 71.4 | 91.3 | 72.2 | 85.3 | 93.8 | 91.7 | 68.9 | 55.6 | 67.7 | 47.8 | 60.0 | 88.4 |
| InternVLM2-2B | 100.0 | 65.2 | 88.9 | 83.8 | 87.5 | 91.7 | 44.4 | 80.6 | 71.0 | 69.6 | 73.3 | 93.0 |
| InternVLM2-8B | 100.0 | 65.2 | 94.4 | 95.6 | 100.0 | 83.3 | 68.9 | 77.8 | 87.1 | 87.0 | 73.3 | 90.7 |
| InternVLM2-40B | 92.9 | 73.9 | 94.4 | 98.5 | 100.0 | 91.7 | 88.9 | 80.6 | 93.5 | 87.0 | 80.0 | 95.3 |
| XgenMM | 92.9 | 91.3 | 83.3 | 88.2 | 93.8 | 95.8 | 60.0 | 75.0 | 77.4 | 52.2 | 66.7 | 90.7 |
| Qwen2-VL | 92.9 | 65.2 | 88.9 | 95.6 | 93.8 | 87.5 | 82.2 | 72.2 | 83.9 | 26.1 | 46.7 | 95.3 |
| Qwen2.5-VL | 100.0 | 87.0 | 94.4 | 97.1 | 100.0 | 91.7 | 68.9 | 66.7 | 90.3 | 47.8 | 40.0 | 90.7 |
| GeminiPro | 92.9 | 78.3 | 50.0 | 94.1 | 87.5 | 83.3 | 60.0 | 11.1 | 80.6 | 34.8 | 6.7 | 86.0 |
| Gemini1.5Pro | 100.0 | 73.9 | 44.4 | 95.6 | 81.2 | 91.7 | 86.7 | 19.4 | 96.8 | 47.8 | 26.7 | 69.8 |
| GPT4V | 85.7 | 73.9 | 50.0 | 69.1 | 81.2 | 91.7 | 57.8 | 19.4 | 93.5 | 30.4 | 13.3 | 69.8 |
| GPT4o-mini | 92.9 | 87.0 | 50.0 | 60.3 | 81.2 | 87.5 | 71.1 | 16.7 | 90.3 | 26.1 | 20.0 | 48.8 |
| GPT4o | 100.0 | 87.0 | 72.2 | 41.2 | 81.2 | 95.8 | 91.1 | 50.0 | 96.8 | 60.9 | 20.0 | 90.7 |
| #1 | #2 | #3 | #4 | #5 | #6 | #7 | #8 | #9 | #11 | #12 | #17 | |
| Dual Acc. | ||||||||||||
| LLaVA1.5-13b | 78.6 | 0.0 | 16.7 | 76.5 | 12.5 | 54.2 | 6.7 | 11.1 | 16.1 | 0.0 | 6.7 | 44.2 |
| LLaVA-NeXT-13B | 71.4 | 0.0 | 38.9 | 85.3 | 43.8 | 58.3 | 53.3 | 52.8 | 51.6 | 4.3 | 20.0 | 79.1 |
| LLaVA-NeXT-34B | 85.7 | 43.5 | 44.4 | 94.1 | 87.5 | 87.5 | 84.4 | 58.3 | 67.7 | 17.4 | 33.3 | 93.0 |
| LLaVA-OV-0.5B | 0.0 | 0.0 | 0.0 | 5.9 | 0.0 | 8.3 | 0.0 | 0.0 | 3.2 | 0.0 | 0.0 | 9.3 |
| LLaVA-OV-7B | 92.9 | 0.0 | 27.8 | 88.2 | 12.5 | 37.5 | 44.4 | 41.7 | 22.6 | 21.7 | 20.0 | 72.1 |
| CogVLM | 0.0 | 0.0 | 0.0 | 33.8 | 6.2 | 4.2 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 16.3 |
| CogVLM2-19B | 85.7 | 0.0 | 38.9 | 89.7 | 18.8 | 58.3 | 22.2 | 30.6 | 16.1 | 0.0 | 6.7 | 65.1 |
| idefics2-8B | 71.4 | 0.0 | 44.4 | 85.3 | 62.5 | 58.3 | 20.0 | 19.4 | 16.1 | 4.3 | 20.0 | 88.4 |
| idefics3-8B | 85.7 | 0.0 | 27.8 | 57.4 | 25.0 | 45.8 | 46.7 | 33.3 | 16.1 | 0.0 | 20.0 | 81.4 |
| Phi3V | 64.3 | 39.1 | 44.4 | 89.7 | 12.5 | 83.3 | 48.9 | 55.6 | 51.6 | 39.1 | 26.7 | 86.0 |
| Phi3.5V | 50.0 | 26.1 | 55.6 | 80.9 | 31.2 | 75.0 | 57.8 | 47.2 | 54.8 | 4.3 | 20.0 | 79.1 |
| InternVLM2-2B | 0.0 | 0.0 | 5.6 | 32.4 | 0.0 | 8.3 | 11.1 | 25.0 | 22.6 | 4.3 | 6.7 | 7.0 |
| InternVLM2-8B | 85.7 | 52.2 | 27.8 | 83.8 | 6.2 | 58.3 | 57.8 | 52.8 | 58.1 | 43.5 | 6.7 | 86.0 |
| InternVLM2-40B | 78.6 | 69.6 | 83.3 | 98.5 | 56.2 | 87.5 | 80.0 | 80.6 | 87.1 | 65.2 | 26.7 | 86.0 |
| XgenMM | 85.7 | 0.0 | 5.6 | 64.7 | 18.8 | 33.3 | 8.9 | 41.7 | 16.1 | 4.3 | 13.3 | 69.8 |
| Qwen2-VL | 85.7 | 47.8 | 72.2 | 94.1 | 50.0 | 62.5 | 77.8 | 63.9 | 48.4 | 13.0 | 40.0 | 95.3 |
| Qwen2.5-VL | 92.9 | 56.5 | 77.8 | 94.1 | 87.5 | 87.5 | 77.8 | 66.7 | 80.6 | 65.2 | 46.7 | 88.4 |
| GeminiPro | 92.9 | 60.9 | 38.9 | 91.2 | 87.5 | 83.3 | 44.4 | 11.1 | 67.7 | 17.4 | 13.3 | 79.1 |
| Gemini1.5Pro | 92.9 | 56.5 | 38.9 | 95.6 | 81.2 | 83.3 | 88.9 | 19.4 | 93.5 | 39.1 | 26.7 | 53.5 |
| GPT4V | 71.4 | 60.9 | 33.3 | 79.4 | 75.0 | 83.3 | 44.4 | 19.4 | 90.3 | 26.1 | 13.3 | 67.4 |
| GPT4o-mini | 85.7 | 78.3 | 38.9 | 36.8 | 87.5 | 83.3 | 57.8 | 8.3 | 87.1 | 30.4 | 6.7 | 20.9 |
| GPT4o | 100.0 | 78.3 | 61.1 | 39.7 | 81.2 | 95.8 | 86.7 | 38.9 | 93.5 | 39.1 | 33.3 | 76.7 |
| UPD Acc. | ||||||||||||
| LLaVA1.5-13b | 85.7 | 0.0 | 16.7 | 80.9 | 18.8 | 54.2 | 8.9 | 19.4 | 22.6 | 4.3 | 33.3 | 53.5 |
| LLaVA-NeXT-13B | 92.9 | 0.0 | 44.4 | 97.1 | 50.0 | 66.7 | 91.1 | 88.9 | 58.1 | 26.1 | 46.7 | 95.3 |
| LLaVA-NeXT-34B | 100.0 | 52.2 | 61.1 | 100.0 | 93.8 | 95.8 | 100.0 | 100.0 | 90.3 | 69.6 | 100.0 | 100.0 |
| LLaVA-OV-0.5B | 0.0 | 0.0 | 5.6 | 11.8 | 0.0 | 8.3 | 2.2 | 8.3 | 12.9 | 4.3 | 20.0 | 18.6 |
| LLaVA-OV-7B | 92.9 | 4.3 | 27.8 | 92.6 | 12.5 | 37.5 | 53.3 | 66.7 | 22.6 | 34.8 | 46.7 | 74.4 |
| CogVLM | 0.0 | 0.0 | 0.0 | 35.3 | 6.2 | 4.2 | 2.2 | 0.0 | 0.0 | 0.0 | 0.0 | 16.3 |
| CogVLM2-19B | 85.7 | 0.0 | 38.9 | 91.2 | 18.8 | 62.5 | 28.9 | 50.0 | 16.1 | 0.0 | 20.0 | 76.7 |
| idefics2-8B | 78.6 | 0.0 | 50.0 | 88.2 | 62.5 | 62.5 | 22.2 | 36.1 | 19.4 | 13.0 | 40.0 | 100.0 |
| idefics3-8B | 92.9 | 0.0 | 33.3 | 60.3 | 25.0 | 45.8 | 62.2 | 52.8 | 16.1 | 0.0 | 26.7 | 88.4 |
| Phi3V | 85.7 | 52.2 | 61.1 | 95.6 | 31.2 | 91.7 | 97.8 | 77.8 | 58.1 | 73.9 | 46.7 | 97.7 |
| Phi3.5V | 85.7 | 26.1 | 61.1 | 97.1 | 31.2 | 83.3 | 88.9 | 88.9 | 58.1 | 13.0 | 40.0 | 95.3 |
| InternVLM2-2B | 0.0 | 0.0 | 5.6 | 42.6 | 0.0 | 12.5 | 17.8 | 25.0 | 22.6 | 4.3 | 13.3 | 9.3 |
| InternVLM2-8B | 85.7 | 65.2 | 27.8 | 88.2 | 12.5 | 75.0 | 82.2 | 75.0 | 64.5 | 56.5 | 20.0 | 97.7 |
| InternVLM2-40B | 92.9 | 82.6 | 94.4 | 100.0 | 56.2 | 95.8 | 97.8 | 100.0 | 100.0 | 87.0 | 46.7 | 97.7 |
| XgenMM | 92.9 | 0.0 | 5.6 | 67.6 | 18.8 | 33.3 | 8.9 | 52.8 | 16.1 | 4.3 | 20.0 | 72.1 |
| Qwen2-VL | 92.9 | 60.9 | 77.8 | 100.0 | 56.2 | 75.0 | 100.0 | 97.2 | 58.1 | 34.8 | 60.0 | 100.0 |
| Qwen2.5-VL | 92.9 | 82.6 | 83.3 | 100.0 | 87.5 | 95.8 | 100.0 | 100.0 | 83.9 | 91.3 | 100.0 | 100.0 |
| GeminiPro | 100.0 | 95.7 | 100.0 | 100.0 | 93.8 | 100.0 | 100.0 | 100.0 | 96.8 | 82.6 | 86.7 | 100.0 |
| Gemini1.5Pro | 100.0 | 95.7 | 100.0 | 100.0 | 87.5 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 97.7 |
| GPT4V | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 |
| GPT4o-mini | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 |
| GPT4o | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 | 96.8 | 100.0 | 100.0 | 100.0 |
| Standard Acc. | ||||||||||||
| LLaVA1.5-13b | 92.9 | 82.6 | 77.8 | 88.2 | 62.5 | 91.7 | 48.9 | 58.3 | 51.6 | 8.7 | 20.0 | 90.7 |
| LLaVA-NeXT-13B | 78.6 | 87.0 | 66.7 | 86.8 | 75.0 | 87.5 | 55.6 | 58.3 | 61.3 | 17.4 | 33.3 | 83.7 |
| LLaVA-NeXT-34B | 85.7 | 82.6 | 77.8 | 94.1 | 93.8 | 91.7 | 84.4 | 58.3 | 71.0 | 39.1 | 33.3 | 93.0 |
| LLaVA-OV-0.5B | 92.9 | 8.7 | 38.9 | 66.2 | 56.2 | 79.2 | 33.3 | 16.7 | 35.5 | 0.0 | 13.3 | 65.1 |
| LLaVA-OV-7B | 100.0 | 47.8 | 94.4 | 95.6 | 100.0 | 95.8 | 82.2 | 69.4 | 87.1 | 69.6 | 73.3 | 93.0 |
| CogVLM | 85.7 | 56.5 | 44.4 | 88.2 | 81.2 | 91.7 | 57.8 | 30.6 | 71.0 | 8.7 | 13.3 | 93.0 |
| CogVLM2-19B | 100.0 | 95.7 | 77.8 | 97.1 | 93.8 | 91.7 | 82.2 | 69.4 | 77.4 | 56.5 | 60.0 | 86.0 |
| idefics2-8B | 92.9 | 78.3 | 83.3 | 95.6 | 81.2 | 91.7 | 73.3 | 36.1 | 54.8 | 52.2 | 60.0 | 88.4 |
| idefics3-8B | 92.9 | 78.3 | 88.9 | 92.6 | 93.8 | 87.5 | 75.6 | 55.6 | 64.5 | 60.9 | 73.3 | 93.0 |
| Phi3V | 78.6 | 82.6 | 83.3 | 92.6 | 81.2 | 91.7 | 51.1 | 66.7 | 74.2 | 52.2 | 73.3 | 88.4 |
| Phi3.5V | 64.3 | 87.0 | 72.2 | 82.4 | 87.5 | 91.7 | 68.9 | 58.3 | 71.0 | 47.8 | 60.0 | 83.7 |
| InternVLM2-2B | 100.0 | 78.3 | 88.9 | 80.9 | 81.2 | 91.7 | 44.4 | 75.0 | 74.2 | 52.2 | 73.3 | 90.7 |
| InternVLM2-8B | 100.0 | 87.0 | 94.4 | 95.6 | 93.8 | 83.3 | 73.3 | 77.8 | 90.3 | 78.3 | 73.3 | 88.4 |
| InternVLM2-40B | 85.7 | 82.6 | 88.9 | 98.5 | 100.0 | 91.7 | 82.2 | 80.6 | 87.1 | 78.3 | 80.0 | 88.4 |
| XgenMM | 92.9 | 87.0 | 83.3 | 91.2 | 93.8 | 95.8 | 60.0 | 75.0 | 77.4 | 56.5 | 73.3 | 90.7 |
| Qwen2-VL | 92.9 | 73.9 | 83.3 | 94.1 | 93.8 | 87.5 | 77.8 | 66.7 | 83.9 | 69.6 | 46.7 | 95.3 |
| Qwen2.5-VL | 100.0 | 73.9 | 94.4 | 94.1 | 100.0 | 91.7 | 77.8 | 66.7 | 96.8 | 73.9 | 46.7 | 88.4 |
| GeminiPro | 92.9 | 60.9 | 38.9 | 91.2 | 87.5 | 83.3 | 44.4 | 11.1 | 67.7 | 30.4 | 13.3 | 79.1 |
| Gemini1.5Pro | 92.9 | 60.9 | 38.9 | 95.6 | 81.2 | 83.3 | 88.9 | 19.4 | 93.5 | 39.1 | 26.7 | 55.8 |
| GPT4V | 71.4 | 60.9 | 33.3 | 79.4 | 75.0 | 83.3 | 44.4 | 19.4 | 90.3 | 26.1 | 13.3 | 67.4 |
| GPT4o-mini | 85.7 | 78.3 | 38.9 | 36.8 | 87.5 | 83.3 | 57.8 | 8.3 | 87.1 | 30.4 | 6.7 | 20.9 |
| GPT4o | 100.0 | 78.3 | 61.1 | 39.7 | 81.2 | 95.8 | 86.7 | 38.9 | 96.8 | 39.1 | 33.3 | 76.7 |